### v4

# Mordecai J. Golin        Albert John Lalim Patupat

Abstract
This paper is split into two parts.
The first addresses the problem of finding a minimum-cost $m$-state Markov chain $(S_{0},\ldots,S_{m-1})$ in a large set of chains. The chains studied have a reward associated with each state. The cost of a chain is its gain, i.e., its average reward under its stationary distribution.
Specifically, for each $k=0,\ldots,m-1,$ there is a known set ${\mathbb{S}}_{k}$ of type-$k$ states. A chain is permissible if it contains exactly one state of each type; the problem is to find a minimum-cost permissible chain.
The original motivation was to find a cheapest binary AIFV-$m$ lossless code on a source alphabet of size $n$. Such a code is an $m$-tuple of trees, in which each tree can be viewed as a Markov chain state. This formulation was then used to address other problems in lossless compression. The known solution techniques for finding minimum-cost Markov chains were iterative and ran in exponential time.
We show how to map type-$k$ states into type-$k$ hyperplanes and then define a Markov Chain Polytope as the lower envelope of all such hyperplanes. A minimum-cost Markov chain can then be shown to be associated with a “highest” point on this polytope.
The local optimization procedures used as subroutines in the previous iterative algorithms are shown to be separation oracles for this polytope. Since these were often polynomial time, an application of the Ellipsoid method for linear programming immediately leads to polynomial time algorithms for these problems.
We illustrate this technique by deriving the first polynomial time algorithms for constructing cheapest binary AIFV-$m$ codes for $m>2.$
1 Introduction
Let $\mathbf{S}=(S_{0},S_{1},\ldots,S_{m-1})$ be an $m$-state Markov chain. For convenience, $[m]$ will denote the set of indices $\{0,1,\ldots,m-1\}.$ See Figure 1.
For $j,k\in[m],$ let $q_{j}(S_{k})$ be the probability of transitioning from state $S_{k}$ to state $S_{j}$. Assume further that $\mathbf{S}$ has a unique stationary distribution
$\bm{\pi}(\mathbf{S})=(\pi_{0}(\mathbf{S}),\ldots,\pi_{m-1}(\mathbf{S})).$
Additionally, each state $S_{k}$ has an associated reward or cost $\ell(S_{k});$ the average steady-state cost or gain of $\mathbf{S}$ is then defined [7] as
${\rm cost}(\mathbf{S})=\sum_{k=0}^{m-1}\ell(S_{k})\pi_{k}(\mathbf{S}).$
A state $S_{k}$ is fully defined by the $m+1$ values $q_{j}(S_{k}),j\in[m]$ and $\ell(S_{k}).$
Figure 1: The bottom half of the figure is a five-state Markov chain. Arrows represent non-zero transition probabilities. $q_{j}(S_{k})$ is the probability of transitioning from state $S_{k}$ to state $S_{j}.$ Note that $q_{0}(S_{k})>0$ for all $k\in[5].$ The circles in the shaded rectangle ${\mathbb{S}}_{k}$ represent the set of all permissible type-$k$ states. Each $S_{k}$ in the Markov Chain shown also belongs to the associated set ${\mathbb{S}}_{k}.$
Next suppose that, for each $k\in[m],$ instead of there being only one $S_{k},$ there exists a large set ${\mathbb{S}}_{k}$ containing all permissible “type-$k$’ states. Fix
${\mathbb{S}}=\bigtimes_{k=0}^{m-1}{\mathbb{S}}_{k}=\left\{(S_{0},S_{1},\ldots,S_{m-1})\mid\forall k\in[m],S_{k}\in{\mathbb{S}}_{k}\right\}$
to be their Cartesian product, and further assume that $\forall\mathbf{S}\in{\mathbb{S}},$ $\mathbf{S}$ has a unique stationary distribition. The problem is to find the Markov chain in ${\mathbb{S}}$ with smallest cost.
This problem first arose in the context of binary AIFV-$m$ coding [4, 22, 23] (described later in detail in Section 6), in which a code is an $m$-tuple $\mathbf{T}=(T_{0},T_{1},\ldots,T_{m-1})$ of binary coding trees for a size $n$ source alphabet; for each $k\in[m],$ there are different restrictions on the structure of $T_{k}.$ The cost of $\mathbf{T}$ is the cost of a corresponding $m$-state Markov chain, so the problem of finding the minimum-cost binary AIFV-$m$ code reduced to finding a minimum-cost Markov chain [5]. This same minimum-cost Markov chain approach was later used to find better parsing trees [17], lossless codes for finite channel coding [18] and AIFV codes for unequal bit-cost coding [15].
Note that in all of these problems, the input size $n$ was relatively small, e.g., a set of $n$ probabilities, but the associated ${\mathbb{S}}_{k}$ had size exponential in $n.$ The previous algorithms developed for solving the problem were iterative ones that moved from Markov chain to Markov chain in ${\mathbb{S}}$, in some non-increasing cost order. For the specific applications mentioned, they ran in exponential (in $n$) time. Each iteration step also required solving a local optimization procedure which was often polynomial time (in $n$).
[8, 10] developed a different approach for solving the binary AIFV-$2$ coding problem, corresponding to a $2$-state Markov chain, in weakly polynomial time using a simple binary search. In those papers, they noted that they could alternatively solve the problem in weakly polynomial time via the Ellipsoid algorithm for linear programming [12] on a two-dimensional polygon. They hypothesized that this latter technique could be extended to $m>2$ but only with a better understanding of the geometry of the problem.
That is the approach followed in this paper, in which we define a mapping of type-$k$ states to type-$k$ hyperplanes in $\mathcal{R}^{m}$. We show that the unique intersection of any $m$ hyperplanes, where each is of a different type, always exists. We call such an intersection point “distinctly-typed” and prove that its “height” is equal to the cost of its associated Markov chain. The solution to the minimum-cost Markov-chain problem is thus the lowest height of any “distinctly-typed’ intersection point.
We then define the Markov-Chain polytope $\mathbb{H}$ to be the lower envelope of the hyperplanes associated with all possible states and note that some lowest-height distinctly-typed intersection point is a highest point on $\mathbb{H}.$ This transforms the problem of finding the cheapest Markov chain to the linear programming one of finding a highest point of $\mathbb{H}.$
The construction and observations described above will be valid for ALL Markov chain problems. In the applications mentioned earlier, the polytope $\mathbb{H}$ is defined by an exponential number of constraints. But, observed from the proper perspective, the local optimization procedures used at each step of the iterative algorithms in [4, 22, 23, 5, 17] can be repurposed as polynomial time separation oracles for $\mathbb{H}$. This permits using the Ellipsoid algorithm approach of [12] to solve the binary AIFV-$m$ problem in weakly polynomial time instead of exponential time.
The remainder of the paper is divided into two distinct parts. Part 1 consists of Sections 2-5. These develop a procedure for solving the generic minimum-cost Markov Chain problem. Section 2 discusses how to map the problem into a linear programming one and how to interpret the old iterative algorithms from this perspective. Section 3 states our new results while Section 4 discusses their algorithmic implications. In particular, Lemma 4.6 states sufficient conditions on ${\mathbb{S}}$ that guarantees a polynomial time algorithm for finding the minimum cost Markov chain. Section 5 then completes Part 1 by proving the main results stated in Section 3.
Part 2, in Sections 6-8, then discusses how to apply Part 1’s techniques to construct best binary AIFV-$m$ codes in weakly polynomial time.
Section 6 provides necessary background, defining binary AIFV-$m$ codes and deriving their important properties. Section 7 describes how to apply the techniques from Section 4 to binary AIFV-$m$ coding. Section 8 proves a very technical lemma specific to binary AIFV-$m$ coding required to show that its associated Markov Chain polytope has a polynomial time separation oracle, which is the last piece needed to apply the Ellipsoid method.
Finally, Section 9 concludes with a quick discussion of other applications of the technique and possible directions for going further.
2 Markov Chains
2.1 The Minimum-Cost Markov Chain problem
Definition 2.1.
Fix $m>1.$
  • (a)
    A state $S$ is defined by a set of $m$ transition probabilities $\{q_{j}(S)\}_{j\in[m]}$ along with a cost $\ell(S).$ $\forall{j\in[m]},\,q_{j}(S)\geq 0$ and $\sum_{j\in[m]}q_{j}(S)=1.$
  • (b)
    $\forall k\in[m],\,$ let ${\mathbb{S}}_{k}$ be some finite given set of states, satisfying that $\forall S_{k}\in{\mathbb{S}}_{k},\,q_{0}(S_{k})>0.$ The states in ${\mathbb{S}}_{k}$ are known as type-$k$ states.
  • (c)
    Markov Chain $\mathbf{S}=(S_{0},\ldots,S_{m-1})$ is permissible if $\forall k\in[m],\,S_{k}\in{\mathbb{S}}_{k}.$
  • (d)
    Define ${\mathbb{S}}=\bigtimes_{k=0}^{m-1}{\mathbb{S}}_{k}=\left\{(S_{0},\ldots,S_{m-1})\mid\forall k\in[m],\,S_{k}\in{\mathbb{S}}_{k}\right\}$ to be the set of permissible Markov chains.
The actual composition and structure of each ${\mathbb{S}}_{k}$ is different from problem to problem and, within a fixed problem, different for different $k.$ The only universal constraint is that stated in (b), that $\forall k\in[m],\,\forall S_{k}\in{\mathbb{S}}_{k},$ $q_{0}(S_{k})>0.$ This implies that $\mathbf{S}=(S_{0},\ldots,S_{m-1})$ is an ergodic unichain, with one aperiodic recurrent class (containing $S_{0}$) and, possibly, some transient states. $\mathbf{S}$ therefore has a unique stationary distribution, (where $\pi_{k}(\mathbf{S})=0$ if and only if $S_{k}$ is a transient state.)
Definition 2.2.
Let $\mathbf{S}=(S_{0},S_{1},\ldots,S_{m-1})$ be a permissible $m$-state Markov chain. The average steady-state cost of $\mathbf{S}$ is defined to be
${\rm cost}(\mathbf{S})=\sum_{k=0}^{m-1}\ell(S_{k})\pi_{k}(\mathbf{S}).$
What has been described above is a Markov Chain with rewards, with ${\rm cost}(\mathbf{S})$ being its gain [7].
Definition 2.3.
The minimum-cost Markov chain problem is to find $\mathbf{S}\in{\mathbb{S}}$ satisfying
${\rm cost}(\mathbf{S})=\min_{\mathbf{S}^{\prime}\in{\mathbb{S}}}{\rm cost}(\mathbf{S}^{\prime}).$
Comments:
(i) In the applications motivating this problem, each ${\mathbb{S}}_{k}$ has size exponential in $n,$ so the search space ${\mathbb{S}}$ has size exponential in $mn.$
(ii) The requirement in (b) that $\forall k\in[m],q_{0}(S_{k})>0,$ (which is satisfied by all the motivating applications,) guarantees that $\mathbf{S}$ has a unique stationary distribution. It is also needed later in other places in the analysis. Whether this condition is necessary is unknown; this is discussed in Section 9.
2.2 Associated Hyperplanes and Polytopes
The next set of definitions map type-$k$ states into type-$k$ hyperplanes in $\mathcal{R}^{m}$ and then defines lower envelopes of those hyperplanes. In what follows, $\mathbf{x}$ denotes a vector $\mathbf{x}=(x_{1},\ldots,x_{m-1})\in\mathcal{R}^{m-1};$ $(\mathbf{x},y)\in\mathcal{R}^{m}$ is a shorthand denoting that $\mathbf{x}\in\mathcal{R}^{m-1}$ and $y\in\mathcal{R}.$ $S_{k}$ denotes a state $S_{k}\in{\mathbb{S}}_{k}.$
Definition 2.4.
Let $k\in[m]$.
  • Define the type-$k$ hyperplanes $f_{k}:\mathcal{R}^{m-1}\times{\mathbb{S}}_{k}\to\mathcal{R}$ as follows:
    $f_{0}(\mathbf{x},S_{0})$ $=$ $\ell(S_{0})+\sum_{j=1}^{m-1}q_{j}(S_{0})\cdot x_{j},$
    $\forall k>0,\ f_{k}(\mathbf{x},S_{k})$ $=$ $\ell(S_{k})+\sum_{j=1}^{m-1}q_{j}(S_{k})\cdot x_{j}-x_{k}.$
  • For all $\mathbf{x}\in\mathcal{R}^{m-1},$ define $g_{k}:\mathcal{R}^{m-1}\to\mathcal{R}$ and $S_{k}:\mathcal{R}^{m-1}\to{\mathbb{S}}_{k}$
    $g_{k}(\mathbf{x})$ $=$ $\min_{S_{k}\in{\mathbb{S}}_{k}}f_{k}(\mathbf{x},S_{k}),$
    $S_{k}(\mathbf{x})$ $=$ $\arg\min_{S_{k}\in{\mathbb{S}}_{k}}f_{k}(\mathbf{x},S_{k}),$
    and Markov chain
    $\mathbf{S}(\mathbf{x})=\left(S_{0}(\mathbf{x}),S_{1}(\mathbf{x})\ldots,S_{m-1}(\mathbf{x})\right).$
    For later use, we note that by definition
    $\forall\mathbf{x}\in\mathcal{R}^{m-1},\,\forall k\in[m],\ g_{k}(\mathbf{x})=f_{k}\left(\mathbf{x},\,S_{k}(\mathbf{x})\right).$ (1)
  • Finally, for all $\mathbf{x}\in\mathcal{R}^{m-1},$ define.
    $h(\mathbf{x})=\min_{k\in[m]}g_{k}(\mathbf{x})=\min_{k\in[m]}\min_{S_{k}\in{\mathbb{S}}_{k}}f_{k}(\mathbf{x},S_{k}).$
$f_{k}$ maps a type-$k$ state $S_{k}$ to a type-$k$ hyperplane $f_{k}(\mathbf{x},S_{k})$ in $\mathcal{R}^{m}.$ For fixed $k,$ $g_{k}(\mathbf{x})$ is the lower envelope of all of the type-$k$ hyperplanes $f_{k}(\mathbf{x},S_{k}).$
Each $S_{k}(\mathbf{x})$ maps point $\mathbf{x}\in\mathcal{R}^{m-1}$ to the lowest type-$k$ hyperplane evaluated at $\mathbf{x}.$ $\mathbf{S}(\mathbf{x})$ maps point $\mathbf{x}\in\mathcal{R}^{m-1}$ to the Markov chain $\mathbf{S}(\mathbf{x})\in{\mathbb{S}}.$
$h(\mathbf{x})$ will be the lower envelope of the $g_{k}(\mathbf{x})$. Since both $g_{k}(\mathbf{x})$ and $h(\mathbf{x})$ are lower envelopes of hyperplanes, they are the upper surface of convex polytopes in $\mathcal{R}^{m}.$ This motivates defining the following polytope:
Definition 2.5.
The Markov Chain Polytope in $\mathcal{R}^{m}$ corresponding to ${\mathbb{S}}$ is
=$\mathbb{H}=\left\{(\mathbf{x},y)\in\mathcal{R}^{m}\mid y\leq h(\mathbf{x})\right\}.$
2.3 The Iterative Algorithm
[4, 5, 22, 23] present an iterative algorithm that was first formulated for finding minimum-cost binary AIFV-$m$ codes and then generalized into a procedure for finding minimum-cost Markov Chains.
The algorithm starts with an arbitrary $\mathbf{x}^{(0)}\in\mathcal{R}^{m-1}$ and iterates, at each step constructing an $\mathbf{x}^{(i)}\in\mathcal{R}^{m-1}$ such that ${\rm cost}\left(\mathbf{S}\left(\mathbf{x}^{(i)}\right)\right)\leq{\rm cost}\left(\mathbf{S}\left(\mathbf{x}^{(i-1)}\right)\right).$ The algorithm terminates at step $i$ if $\mathbf{x}^{(i)}=\mathbf{x}^{(i-1)}.$
Although it does not run in polynomial time, it is noteworthy because the fact of its termination (Lemma 2.6) will be needed later (in Corollary 3.4) to prove that $\mathbb{H}$ contains some point corresponding to a minimum-cost Markov chain.
[23] proves that the algorithm always terminates when $m=2$ and, at termination, $\mathbf{S}\left(\mathbf{x}^{(i)}\right)$ is a minimum-cost Markov Chain. [4, 5, 22] prove that [^2] They also claim that the algorithm always terminates. The proofs of correctness there are only sketches and missing details but they all seem to implictly assume that $q_{j}(S_{k})>0$ for all $j,k\in[m]$ and not just $j=0.$ , for $m>2,$ if the algorithm terminates, then $\mathbf{S}\left(\mathbf{x}^{(i)}\right)$ is a minimum-cost Markov Chain. A complete proof of termination (and therefore solution of the minimum-cost Markov Chain problem) is provided in [1]. This states
Lemma 2.6.
[Theorem 1, [1]] No matter what the starting value $\mathbf{x}^{(0)},$ there always exists $i$ such that $\mathbf{x}^{(i)}=\mathbf{x}^{(i-1)}.$ Furthermore, for that $i,$
$g_{0}\left(\mathbf{x}^{(i)}\right)=g_{1}\left(\mathbf{x}^{(i)}\right)=\cdots=g_{m-1}\left(\mathbf{x}^{(i)}\right).$
Notes/Comments:
  • The algorithm in [4, 5, 22, 23] looks different than the one described in [1] but they are actually identical, just expressed in a different coordinate system. The relationship between the two coordinate systems is shown in [11]. The coordinate system in [1] is the same as the one used in this paper.
  • Each step of the iterative algorithm requires calculating $S_{k}(\mathbf{x}^{(t)})$ for the current $t$ and all $k\in[m].$ In applications, finding $S_{k}(\mathbf{x})$ is very problem specific and is usually a combinatorial optimization problem. For example, the first papers on AIFV-$m$ coding [23] and the most current papers on finite-state channel coding [18] calculate them using integer linear programming, as does a recent paper on constructing AIFV codes for unequal bit costs [15]. The more recent papers on both AIFV-$m$ coding [16, 22, 11] and AIVF coding [17] use dynamic programming.
  • As noted, the fact that if the iterative algorithm terminates at some $\mathbf{x}^{(i)}$ then $\mathbf{S}\left(\mathbf{x}^{(i)}\right)$ is a minimum-cost Markov Chain was proven directly in [4, 5, 22, 23]. An alternative proof of this fact is given in Lemma 3.3 in this paper.
  • The proof of termination given in [1] Theorem 1, strongly depends upon condition (b) from Definition 2.1, i.e., $\forall k\in[m],\,q_{0}(S_{k})>0.$
3 The Main Results
This section states our two main lemmas, Lemma 3.1 and Lemma 3.7, and their consequences. Their proofs are deferred to Section 5.
$\begin{array}[]{|l||l|l|l|l||l|}\hline\cr i&q_{0}(S_{i})&q_{1}(S_{i})&q_{2}(S_{i})&\ell(S_{i})\\ \hline\cr\hline\cr 0&0.5&0.25&0.25&9\\ \hline\cr 1&0.75&0&0.25&11\\ \hline\cr 2&0.75&0.25&0&14\\ \hline\cr\end{array}$
Figure 2: An illustration of Lemma 3.1 (a) for a 3-state Markov chain $\mathbf{S}=(S_{0},S_{1},S_{2}).$ The table lists the associated $q_{j}(S_{k})$ and $\ell(S_{k})$ values. The green plane is $f_{0}(\mathbf{x},S_{0})=9+x_{1}/4+x_{2}/4,$ the red plane $f_{1}(\mathbf{x},S_{1})=11-x_{1}+x_{2}/4,$ and the blue plane $f_{2}(\mathbf{x},S_{2})=14+x_{1}/4-x_{2}.$ By calculation, $\bm{\pi}(\mathbf{S})=(0.6,0.15,0.25)$ so $\mbox{\rm cost}(\mathbf{S})=0.6\cdot 9+0.15\cdot 11+0.25\cdot 14=10.55.$ The planes intersect at unique point $(x_{0},x_{1},y)=(1.6,4.6,10.55)$.
Lemma 3.1 (Geometric Properties).
Let $\mathbf{S}=(S_{0},\ldots,S_{m-1})\in{\mathbb{S}}$ be a permissible Markov Chain.
  • (a)
    The $m$ $(m-1)$-dimensional hyperplanes $y=f_{k}(\mathbf{x},S_{k}),$ $k\in[m],$ intersect at a unique point $(\mathbf{x}^{int},y^{int})\in\mathcal{R}^{m}$.
    We call such a $(\mathbf{x}^{int},y^{int})$ point a distinctly-typed intersection point.
  • (b)
    $\forall\mathbf{x}\in\mathcal{R}^{m-1},\quad y^{int}={\rm cost}(\mathbf{S})=\sum_{k=0}^{m-1}f_{k}(\mathbf{x},S_{k})\cdot\pi_{k}(\mathbf{S}).$
  • (c)
    The intersection point $(\mathbf{x}^{int},y^{int})$ is on or above the lower envelope of the $m$ hyperplanes $f_{k}(\mathbf{x},S_{k}),$ $k\in[m],$ i.e.,
    $\forall\ \mathbf{x}\in\mathcal{R}^{m-1},\quad y^{int}\geq\min_{k\in[m]}f_{k}(\mathbf{x},S_{k}).$
  • (d)
    The intersection point $(\mathbf{x}^{int},y^{int})$ also satisfies
    $\forall\ \mathbf{x}\in\mathcal{R}^{m-1},\quad y^{int}\geq h(\mathbf{x}).$
The Lemma immediately implies that finding a minimum-cost Markov Chain is equivalent to finding a minimum “height” ($m$’th-coordinate) distinctly-typed intersection point. It also implies the following corollary.
Corollary 3.2.
Let $\mathbf{S}=(S_{0},\ldots,S_{m-1})\in{\mathbb{S}}$.
$\min_{\mathbf{S}\in{\mathbb{S}}}{\rm cost}(\mathbf{S})\geq\max_{\mathbf{x}\in\mathcal{R}^{m-1}}h(\mathbf{x}).$
This in turn, permits, proving
Lemma 3.3.
If
$g_{0}(\mathbf{x}^{*})=g_{1}(\mathbf{x}^{*})=\cdots=g_{m-1}(\mathbf{x}^{*})=y^{*}$ (2)
for some $\mathbf{x}^{*}\in\mathcal{R}^{m-1}$ and $y^{*}\in\mathcal{R},$ then
$y^{*}=h(\mathbf{x}^{*})={\rm cost}\left(\mathbf{S}\left(\mathbf{x}^{*}\right)\right)$ (3)
and
$\min_{\mathbf{S}\in{\mathbb{S}}}{\rm cost}(\mathbf{S})={\rm cost}(\mathbf{S}(\mathbf{x}^{*}))=h(\mathbf{x}^{*})=\max_{\mathbf{x}\in\mathcal{R}^{m-1}}h(\mathbf{x})=\max\left\{y\mid(\mathbf{x},y)\in\mathbb{H}\right\}.$ (4)
Proof.
By the definition of the $g_{i}$ and $h,$ if $\mathbf{x}^{*}$ satisfies (2), $h\left(\mathbf{x}^{*}\right)=y^{*}.$
Recall that for all $k\in[m],$ $g_{k}(\mathbf{x})=f_{k}(\mathbf{x},S_{k}(\mathbf{x})).$ Thus, (2), implies that the hyperplanes $f_{k}(\mathbf{x},S_{k}(\mathbf{x}^{*}))$ intersect at the point $\left(\mathbf{x}^{*},y^{*}\right)=\left(\mathbf{x}^{*},h\left(\mathbf{x}^{*}\right)\right).$ Applying Lemma 3.1 (b) proves (3).
Combining (3) with Corollary 3.2 immediately implies
$h(\mathbf{x}^{*})={\rm cost}(\mathbf{S}(\mathbf{x}^{*}))\geq\min_{\mathbf{S}\in{\mathbb{S}}}{\rm cost}(\mathbf{S})\geq\max_{\mathbf{x}\in\mathcal{R}^{m-1}}h(\mathbf{x})\geq h(\mathbf{x}^{*}).$ (5)
Because the leftmost and rightmost values in (5) are identical, the two inequalities in (5) must be equalities, proving (4). ∎
Condition (2) in Lemma 3.3 means that the $m$ different lower envelopes $g_{k}(\mathbf{x})$, $k\in[m],$ must simultaneously intersect at a point $\mathbf{x}^{*}$. It is not a-priori obvious that such an $\mathbf{x}^{*}$ should always exist but Lemma 2.6 tells us that the iterative algorithm always terminates at such an $\mathbf{x}^{*}$, immediately proving [^3] For completness, we note that Lemma 7.4 later also directly proves that condition (2) in Lemma 3.3 holds for the special case of AIFV-$m$ coding. Thus this paper provides a fully self-contained proof of Corollary 3.4 for AIFV-$m$ coding.
Corollary 3.4.
There exists $\mathbf{x}^{*}\in\mathcal{R}^{m-1}$ satisfying (2) in Lemma 3.3. This $\mathbf{x}^{*}$ satisfies $\left(\mathbf{x}^{*},h\left(\mathbf{x}^{*}\right)\right)\in\mathbb{H}$ and $h\left(\mathbf{x}^{*}\right)=\max\left\{y\mid(\mathbf{x},y)\in\mathbb{H}\right\}.$
Figure 3: An illustration of Lemma 3.3 and Corollary 3.4 for the case $m=3.$ 3000 states each of type 00, $1$ and $2$ were generated with associated $\ell(S_{k})$ values. The green, red and blue are, respectively the lower envelopes $g_{0}(\mathbf{x})$, $g_{1}(\mathbf{x})$, $g_{2}(\mathbf{x}),$ i.e., the lower envelopes of the 3000 associated hyperplanes of each type. $\mathbb{H}$ is the lower envelope of those three lower envelopes. The three $g_{i}(\mathbf{x})$ intersect at a unique point $(x_{1}^{*},x_{2}^{*},y^{*})$ which is a highest point in $\mathbb{H}.$ $\mathbf{S}^{*}=\mathbf{S}((x_{1}^{*},x_{2}^{*}))$ is a minimal cost Markov chain among the $3000^{3}$ permissible Markov chains in ${\mathbb{S}}$ and ${\rm cost}(\mathbf{S}^{*})=y^{*}$.
This suggests a new approach to solving the minimum-cost Markov chain problem.
  1. 1.
    Use linear programming to find a highest point $(\hat{\mathbf{x}},\hat{y})\in\mathbb{H}.$
  2. 2.
    Starting from $\hat{\mathbf{x}},$ find a distinctly-typed intersection point $(\mathbf{x}^{*},\hat{y})\in\mathbb{H}.$
  3. 3.
    Return $\mathbf{S}(\mathbf{x}^{*}).$
From Corollary 3.4, an $\mathbf{x}^{*}$ in step 2 must exist; from Lemma 3.3 any such $\mathbf{x}^{*}$ found would yield a $S(\mathbf{x}^{*}),$ solving the minimum-cost Markov chain problem.
While it would usually not be difficult to move to a highest vertex $\mathbf{x}^{*}\in\mathbb{H}$ when starting from a highest point $\hat{\mathbf{x}}\in\mathbb{H},$ the added complication that $\mathbf{x}^{*}$ must be distinctly-typed complicates step 2. These complications can be sidestepped using the extension to $P$-restricted search spaces and the pruning procedure described below.
Definition 3.5 ($P$-restricted search spaces).
Let $\mathcal{P}$ denote the set of all subsets of $[m]$ that contain “0”, i.e., $\mathcal{P}=\{P\subseteq[m]\mid 0\in P\}.$ For all $k\in[m]$ and all $s_{k}\in{\mathbb{S}}_{k},$ define $P(S_{k})=\{j\in[m]\,:\,q_{j}(S_{k})>0\},$ the set of all states to which $S_{k}$ can transition. Since $q_{0}(S_{k})>0,$ $P(S_{k})\in\mathcal{P}.$
Now fix $k\in[m]$ and $P\in\mathcal{P}.$ Define
${\mathbb{S}}_{k|P}=\Bigl{\{}S_{k}\in{\mathbb{S}}_{k}\mid P(S_{k})\subseteq P\Bigr{\}}.$
to be the subset of states in ${\mathbb{S}}_{k}$ that only transition to states in $P.$
Further define
${\mathbb{S}}_{|P}=\bigtimes_{k=0}^{m-1}{\mathbb{S}}_{k|P}.$
Note that ${\mathbb{S}}_{k|[m]}={\mathbb{S}}_{k}$ and ${\mathbb{S}}_{|[m]}={\mathbb{S}}$. Also note that if $(S_{0},\ldots,S_{m-1})=\mathbf{S}\in{\mathbb{S}}_{|P}$ then, if $k\not\in P,$ $S_{k}$ is a transient state in $\mathbf{S}.$
Definition 3.6.
Let $k\in[m]$ and $P\in\mathcal{P}.$
For all $\mathbf{x}\in\mathcal{R}^{m-1},$ define $g_{k|P}:\mathcal{R}^{m-1}\to\mathcal{R}$ and $S_{k|P}:\mathcal{R}^{m-1}\to{\mathbb{S}}_{k|P}$ as
$g_{k|P}(\mathbf{x})$ $=$ $\min_{S_{k}\in{\mathbb{S}}_{k|P}}f_{k}(\mathbf{x},S_{k}),$
$S_{k|P}(\mathbf{x})$ $=$ $\arg\min_{S_{k}\in{\mathbb{S}}_{k|P}}f_{k}(\mathbf{x},S_{k}),$
and
$\mathbf{S}_{|P}(\mathbf{x})=\left(S_{0|P}(\mathbf{x}),S_{1|P}(\mathbf{x}),\ldots,S_{m-1|P}(\mathbf{x})\right).$
Note that $g_{k}(\mathbf{x})$, $S_{k}(\mathbf{x})$ and $\mathbf{S}(\mathbf{x})$ are, respectively, equivalent to $g_{k|[m]}(\mathbf{x})$, $S_{k|[m]}(\mathbf{x})$ and $\mathbf{S}_{|[m]}(\mathbf{x}),$ i.e, all types permitted.
These definitions permit moving from any “highest” point on $\mathbb{H}$ to a distinctly-typed intersection point at the same height.
Lemma 3.7.
Suppose $(\hat{\mathbf{x}},\hat{y})\in\mathbb{H}$ satisfies $\hat{y}=\max\left\{y\,:\,(\mathbf{x},y)\in\mathbb{H}\right\}.$
Let $\mathbf{S}^{*}=(S^{*}_{0},S^{*}_{1},\dots,S^{*}_{m-1})$ denote any minimum-cost Markov chain and $P^{*}=\{k:\pi_{k}(\mathbf{S}^{*})>0\},$ the set of its recurrent indices $k$. Let $P\in\mathcal{P}.$ Then
  • (a)
    If $\forall k\in P,$ $g_{k|P}(\hat{\mathbf{x}})=\hat{y}$, then ${\rm cost}\left(\mathbf{S}_{|P}(\hat{\mathbf{x}})\right)=\hat{y}.$
    Equivalently, this implies $\mathbf{S}_{|P}(\hat{\mathbf{x}})$ is a minimum-cost Markov chain.
  • (b)
    If $P^{*}\subseteq P$, then $\{0\}\subseteq P^{*}\subseteq\{k\in P:g_{k|P}(\hat{\mathbf{x}})=\hat{y}\}$.
Corollary 3.8 (Iterative Pruning Yields Optimal Solution).
Suppose $(\hat{\mathbf{x}},\hat{y})\in\mathbb{H}$ satisfies $\hat{y}=\max\left\{y\,:\,(\mathbf{x},y)\in{\mathbb{H}}\right\}.$
Then the procedure $\mbox{\rm Prune}(\mathbf{x})$ in Algorithm 6 terminates in at most $m-1$ steps. At termination, $\mathbf{S}_{|R}(\hat{\mathbf{x}})$ is a minimum-cost Markov chain.
1:Set $R=[m]$
2:Set $R^{\prime}=\{k\in[m]\mid g_{k|[m]}(\hat{\mathbf{x}})=\hat{y}\}.$
3:while $R^{\prime}\not=R$ do
4:     Set $R=R^{\prime}$
5:     Set $R^{\prime}=\{k\in R\mid g_{k|R}(\hat{\mathbf{x}})=\hat{y}\}.$
6:end while
Algorithm 1 $\mbox{Prune}(\hat{\mathbf{x}}):$ $(\hat{\mathbf{x}},\hat{y})\in{\mathbb{H}}$ where $\hat{y}=\max\left\{y\,:\,(\mathbf{x},y)\in{\mathbb{H}}\right\}.$
Proof.
At the start of the algorithm, $P^{*}\subseteq[m]=R.$ As long as $R^{\prime}\not=R$, Lemma 3.7 (b) implies that $P^{*}\subseteq R^{\prime}\subseteq R.$ If $R\not=R^{\prime}$, the size of $R$ decreases, so the process can not run more than $m-1$ steps. At termination, $R^{\prime}=R,$ so from Lemma 3.7 (a), $\mathbf{S}_{|R}(\hat{\mathbf{x}})$ is a minimum-cost Markov chain. ∎
4 Algorithmic Implications
Stating our running times requires introducing further definitions.
Definition 4.1.
Let ${\bf R}=\bigtimes_{k\in[m]}[\ell_{k},r_{k}]$ be a hyperrectangle in $\mathcal{R}^{m-1}.$ Define
  • $t_{{\mathbb{S}}}({\bf R})$ is the maximum time required to calculate $\mathbf{S}(\mathbf{x})$ for any $\mathbf{x}\in{\bf R}.$
  • $t^{\prime}_{{\mathbb{S}}}({\bf R})$ is the maximum time required to calculate $\mathbf{S}_{|P}(\mathbf{x})$ for any $\mathbf{x}\in{\bf R}$ and any $P\in\mathcal{P}.$
Note that for any ${\bf R},$ $t_{{\mathbb{S}}}({\bf R})\leq t^{\prime}_{{\mathbb{S}}}({\bf R}).$ Also, $t_{{\mathbb{S}}}$ and $t^{\prime}_{{\mathbb{S}}}$ will denote $t_{{\mathbb{S}}}\left(\mathcal{R}^{m-1}\right)$ and $t^{\prime}_{{\mathbb{S}}}\left(\mathcal{R}^{m-1}\right).$
In addition,
  • $\mathcal{I}$ denotes the number of iterative steps made by the iterative algorithm. The only known bound for this is $|{\mathcal{I}}|\leq|{\mathbb{S}}|,$ the number of permissible Markov Chains.
The iterative algorithm requires $O({\mathcal{I}}\cdot t_{{\mathbb{S}}})$ time. Improvements to its running time have focused on improving $t_{{\mathbb{S}}}$ in specific applications.
For binary AIFV-$m$ coding, of $n$ source-code words, $\mathbf{S}(\mathbf{x})$ was first solved using integer linear programming [23] so $t_{{\mathbb{S}}}$ was exponential in $n.$ For the specific case of ${\bf R}=[0,1]^{m-1},$ this was improved to polynomial time using different dynamic programs. More specifically, for $m=2,$ [16] showed that $t_{{\mathbb{S}}}({\bf R})=O(n^{5})$, improved to $O(n^{3})$ by [9]; for $m>2,$ [22] showed that $t_{{\mathbb{S}}}({\bf R})=O(n^{2m+1})$, improved to $O(n^{m+2})$ by [11]. These sped up the running time of the iterative algorithm under the (unproven) assumption that the iterative algorithm always stayed within $[0,1]^{m-1}$.
For AIVF codes, [17] proposed using a modification of a dynamic programming algorithm due to [3], yielding that for any fixed $m$, $t_{{\mathbb{S}}}$ is polynomial time in $D,$ the number of words permitted in the parse dictionary.
In all of these cases, though, the running time of the algorithm is still exponential because ${\mathcal{I}}$ could be exponential.
Note: Although $t_{{\mathbb{S}}}({\bf R})$ has been studied, nothing was previously known about $t^{\prime}_{{\mathbb{S}}}({\bf R}).$ This is simply because there was no previous need to define and construct $\mathbf{S}_{|P}(\mathbf{x})$. In all the known cases in which algorithms for constructing $\mathbf{S}(\mathbf{x})$ exist, it is easy to slightly modify them to construct $\mathbf{S}_{|P}(\mathbf{x})$ in the same running time. So $t^{\prime}_{{\mathbb{S}}}({\bf R})=\Theta(t_{{\mathbb{S}}}({\bf R})).$
We now see how the properties of the Markov chain polytope will, under some fairly loose conditions, permit finding the minimum-cost Markov chain in polynomial time. This will be done via the Ellipsoid method of Grötschel, Lovász and Schrijver [12, 13, 20], which, given a polynomial time separation oracle for a polytope, permits solving a linear programming problem on the polytope in polynomial time. The main observation is that $\mathbf{S}(\mathbf{x})$ provides a separation oracle for $\mathbb{H},$ so the previous application specific algorithms for constructing $\mathbf{S}(\mathbf{x})$ can be reused to derive polynomial time algorithms.
4.1 Separation oracles and $\mathbb{H}$
Recall the definition of a separation oracle.
Definition 4.2 ([20]).
Let $K\subseteq\mathbb{R}^{m}$ be a closed convex set. A separation oracle [^4] Some references label this a strong separation oracle. We follow the formulation of [20] in not adding the word strong. for $K$ is a procedure that, for any $\mathbf{x}\in\mathbb{R}^{m},$ either reports that $\mathbf{x}\in K$ or, if $\mathbf{x}\not\in K$, returns a hyperplane that separates $\mathbf{x}$ from $K.$ That is, it returns $\mathbf{a}\in\mathbb{R}^{m}$ such that $\forall\mathbf{z}\in K,$ ${\mathbf{a}}^{T}\mathbf{x}>\mathbf{a}^{T}\mathbf{z}.$
It is now clear that $\mathbf{S}(\mathbf{x})$ provides a separation oracle for $\mathbb{H}.$
Lemma 4.3.
Let $m$ be fixed and $\mathbb{H}$ be the Markov Chain polytope. Let $\mathbf{z}=(\mathbf{x},y)\in\mathcal{R}^{m}$. Then knowing $\mathbf{S}(\mathbf{x})=(S_{0}(\mathbf{x}),\dots,S_{m-1}(\mathbf{x}))$ provides a $O(m^{2})$ time algorithm for either reporting that $\mathbf{z}\in\mathbb{H}$ or returning a hyperplane that separates $\mathbf{z}$ from $\mathbb{H}.$
Proof.
$\mathbf{z}\in\mathbb{H}$ if and only if
$y\leq h(\mathbf{x})=\min_{k\in[m]}\left(\min_{S_{k}\in{\mathbb{S}}_{k}}f_{k}(\mathbf{x},S_{k})\right)=\min_{k\in[m]}g_{k}(\mathbf{x})=\min_{k\in[m]}f_{k}\left(\mathbf{x},S_{k}(\mathbf{x})\right).$
Thus knowing $\mathbf{S}(\mathbf{x})$ immediately determines whether $\mathbf{z}\in\mathbb{H}$ or not. Furthermore if $\mathbf{z}\not\in\mathbb{H}$, i.e., $y^{\prime}>h(\mathbf{x}),$ let $k^{\prime}\in[m]$ be an index satisfying $f_{k^{\prime}}\left(\mathbf{x},S_{k^{\prime}}(\mathbf{x})\right)=h(\mathbf{x}).$
The hyperplane $y=f_{k^{\prime}}(\mathbf{x},S_{k^{\prime}}(\mathbf{x}))$ then separates $(\mathbf{x},y^{\prime})$ from $\mathbb{H}$ because $y=f_{k^{\prime}}(\mathbf{x},S_{k^{\prime}}(\mathbf{x}))$ is a supporting hyperplane of $\mathbb{H}$ at point $(\mathbf{x},h(\mathbf{x})).$
4.2 The Ellipsoid Algorithm with Separation Oracles
The ellipsoid method of Grötschel, Lovász and Schrijver [12, 13] states that, given a polynomial-time separation oracle, (even if $K$ is a polytope defined by an exponential number of hyperplanes) an approximate optimal solution to the “convex optimization problem” can be found in polynomial time. If $K$ is a rational polytope, then an exact optimal solution can be found in polynomial time. We follow the formulation of [20] in stating these results.
Definition 4.4 ([20] Section 14.2).
[…]
  • (i)
    give a vector $\mathbf{x}_{0}\in K$ with $\mathbf{c}^{T}\mathbf{x}_{0}=\max\{\mathbf{c}^{T}\mathbf{x}\,|\,\mathbf{x}\in K\}.$
  • (ii)
[…]
  • (iii)
    assert that $K$ is empty.
[…]
The relevant result of [12, 13] is
Theorem 4.5 ([20] Corollary 14.1a).
There exists an algorithm ELL such that if ELL is given the input $(m,\varphi,\mbox{\tt SEP},c)$ where:
          $m$ and $\varphi$ are natural numbers and SEP is a separation oracle for some rational polyhedron $K$ in $\mathcal{R}^{m}$, defined by linear inequalities of size at most $\varphi$ and $c\in\mathbb{Q}^{m},$
[…]
In this statement, the size of a linear inequality is the number of bits needed to write the rational coefficients of the inequality, where the number of bits required to write rational $r=p/q,$ where $p,q$ are relatively prime integers, is $\lceil\log_{2}p\rceil+\lceil\log_{2}q\rceil.$
4.3 Solving the Minimum-Cost Markov Chain problem.
Combining all of the pieces, we can now prove our main result.
Lemma 4.6.
Given ${\mathbb{S}},$ let $\varphi$ be the maximum number of bits required to write any transition probability $q_{i}(S)$ or cost $\ell(S)$ of a permissible state $S$.
Furthermore, assume some known hyper-rectangle ${\bf R}\subset\mathcal{R}^{m-1}$ with the property that there exists $(\mathbf{x}^{*},y^{*})\in\mathbb{H}$ satisfying $\mathbf{x}^{*}\in{\bf R}$ and $y^{*}=\max\{y^{\prime}\,:\,(\mathbf{x}^{\prime},y^{\prime})\in\mathbb{H}\}$.
Then the minimum-cost Markov chain problem can be solved in time polynomially bounded by $m$, $\varphi$ and $t^{\prime}_{S}({\bf R}).$
Proof.
Without loss of generality, we assume that $y^{*}\geq 0.$ To justify, recall that the minimum cost Markov chain $\mathbf{S}^{*}=(S^{*}_{0},\ldots,S^{*}_{m-1})$ has cost $y^{*}=\sum_{k}\ell(S^{*}_{k})\pi_{k}(\mathbf{S}^{*})\geq\min_{k}\ell(S^{*}_{k})$ where $\pi_{k}(\mathbf{S}^{*})\geq 0$ is the $k$’th component of the stationary distribution of $\mathbf{S}^{*}.$ Trivially, if $\ell(S^{*}_{k})\geq 0$ for all $k,$ then $y^{*}\geq 0.$
The original problem formulation does not require that $\ell(S_{k})\geq 0$. But, we can modify a given input by adding the same constant $2^{\varphi}$ to $\ell(S_{k})$ for every $S_{k}\in{\mathbb{S}}_{k}$, $k\in[m].$ This makes every $\ell(S_{k})$ non-negative so the minimum cost Markov chain in this modified problem has non-negative cost. Since this modification adds $2^{\varphi}$ to the cost of every Markov chain, solving the modified problem solves the original problem. Note that this modification can at most double $\varphi,$ so this does not break the statement of the lemma. We may thus assume that $y^{*}\geq 0$.
Thus $(\mathbf{x}^{*},y^{*})\in\mathbb{H}^{\prime}=\mathbb{H}\cap{\bf R}^{\prime}$ where ${\bf R}^{\prime}=\left\{(\mathbf{x},y)\,:\,\mathbf{x}\in{\bf R},\,y\geq 0\right\}\subset\mathcal{R}^{m}.$ Since $y^{*}$ is bounded from above by the cost of any permissible Markov chain, $\mathbb{H}^{\prime}$ is a bounded non-empty polytope.
Since $(\mathbf{x}^{*},y^{*})\in\mathbb{H}^{\prime},$ $\max\{y^{\prime}\,:\,(\mathbf{x}^{\prime},y^{\prime})\in\mathbb{H}\}=\max\{y^{\prime}\,:\,(\mathbf{x}^{\prime},y^{\prime})\in\mathbb{H}^{\prime}\}.$
Now consider the following separation oracle for $\mathbb{H}^{\prime}.$ Let $(\mathbf{x},y)\in\mathcal{R}^{m}.$
  • In $O(1)$ time, first check whether $y\geq 0.$ If no, then $(\mathbf{x},y)\not\in\mathbb{H}^{\prime}$ and $y=0$ is a separating hyperplane.
  • Otherwise, in $O(m)$ time, check whether $\mathbf{x}\in{\bf R}.$ If no, $(\mathbf{x},y)\not\in\mathbb{H}^{\prime}$ and a separating hyperplane is just the corresponding side of $\bf R$ that $(\mathbf{x},y)$ is outside of.
  • Otherwise, calculate $\mathbf{S}(\mathbf{x})$ in $t_{S}({\bf R})$ time. From Lemma 4.3, this provides a separation oracle.
Consider solving the optimization problem on polytope $K=\mathbb{H}^{\prime}$ with $\mathbf{c}=(0,0,\ldots,0,1)$ to find $(\mathbf{x}_{0},y_{0})\in\mathbb{H}^{\prime}$ satisfying
$\mathbf{c}^{T}(\mathbf{x}_{0},y_{0})=\max\{\mathbf{c}^{T}(\mathbf{x}^{\prime},y^{\prime})\,:\,(\mathbf{x}^{\prime},y^{\prime})\in\mathbb{H}^{\prime}\}=\max\{y^{\prime}\,:\,(\mathbf{x}^{\prime},y^{\prime})\in\mathbb{H}^{\prime}\}=y^{*}.$
Since $\mathbb{H}^{\prime}$ is a bounded non-empty polytope, we can apply Theorem 4.5 to find such an $(\mathbf{x}_{0},y_{0})$ in time polynomially bounded in $m,$ $\varphi$, and $t_{{\mathbb{S}}}({\bf R}).$
Applying Corollary 3.8 and its procedure $\mbox{\rm Prune}(\mathbf{x}_{0})$ then produces a minimum cost Markov-Chain in another $O(mt^{\prime}_{{\mathbb{S}}}({\bf R}))$ time. The final result follows from the fact that $t_{{\mathbb{S}}}({\bf R})\leq t^{\prime}_{{\mathbb{S}}}({\bf R}).$
Part 2, starting in Section 6, shows how to apply this Lemma to derive a polynomial time algorithm for constructing minimum-cost AIFV-$m$ codes.
We also note that [2] recently applied Lemma 4.3 in a plug-and-play manner to derive the first polynomial time algorithms for constructing optimal AIVF codes. AIVF codes are a multi-tree generalization of Tunstall coding [17, 18], for which the previous constriction algorithms ran in exponential time.
5 Proofs of Lemmas 3.1 and 3.7
5.1 Proof of Lemma 3.1
Before starting the proof, we note that (a) and the first equality of (b) are, after a change of variables, implicit in the analysis provided in [5] of the convergence of their iterative algorithm. The derivation there is different than the one provided below, though, and is missing intermediate steps that we need for proving our later lemmas.
Proof.
In what follows, $\mathbf{Q}$ denotes $\mathbf{Q}(\mathbf{S})$, the transition matrix associated with $\mathbf{S},$ and $\bm{\pi}$ denotes $\bm{\pi}(\mathbf{S})$, its unique stationary distribution.
To prove (a) observe that the intersection condition
$y^{int}=f_{0}(\mathbf{x}^{int},S_{0})=f_{1}(\mathbf{x}^{int},S_{1})=\dots=f_{m-1}(\mathbf{x}^{int},S_{m-1})$ (6)
can be equivalently rewritten as
$\begin{bmatrix}y^{int}\\ y^{int}\\ \vdots\\ y^{int}\\ \end{bmatrix}=\begin{bmatrix}f_{0}(\mathbf{x}^{int},S_{0})\\ f_{1}(\mathbf{x}^{int},S_{1})\\ \vdots\\ f_{m-1}(\mathbf{x}^{int},S_{m-1})\\ \end{bmatrix},$ (7)
where the right-hand side of (7) can be expanded into
$\begin{bmatrix}\ell(S_{0})\\ \ell(S_{1})\\ \vdots\\ \ell(S_{m-1})\\ \end{bmatrix}+\begin{bmatrix}q_{0}(S_{0})&q_{1}(S_{0})&\dots&q_{m-1}(S_{0})\\ q_{0}(S_{1})&q_{1}(S_{1})&\dots&q_{m-1}(S_{1})\\ \vdots&\vdots&\ddots&\vdots\\ q_{0}(S_{m-1})&q_{1}(S_{m-1})&\dots&q_{m-1}(S_{m-1})\\ \end{bmatrix}\begin{bmatrix}0\\ x^{int}_{1}\\ \vdots\\ x^{int}_{m-1}\\ \end{bmatrix}-\begin{bmatrix}0\\ x^{int}_{1}\\ \vdots\\ x^{int}_{m-1}\\ \end{bmatrix}.$
Equation (7) can therefore be rewritten as
$-\begin{bmatrix}\ell(S_{0})\\ \ell(S_{1})\\ \vdots\\ \ell(S_{m-1})\\ \end{bmatrix}=\begin{bmatrix}1&q_{1}(S_{0})&\dots&q_{m-1}(S_{0})\\ 1&q_{1}(S_{1})-1&\dots&q_{m-1}(S_{1})\\ \vdots&\vdots&\ddots&\vdots\\ 1&q_{1}(S_{m-1})&\dots&q_{m-1}(S_{m-1})-1\\ \end{bmatrix}\begin{bmatrix}-y^{int}\\ x^{int}_{1}\\ \vdots\\ x^{int}_{m-1}\\ \end{bmatrix},$ (8)
where the matrix in (8), denoted as $\mathbf{M},$ is $\mathbf{Q}$ after subtracting the identity matrix $\mathbf{I}$ and replacing the first column with $1$s. To prove (a) it therefore suffices to prove that $\mathbf{M}$ is invertible.
The uniqueness of $\bm{\pi}$ implies that the kernel of $\mathbf{Q}-\mathbf{I}$ is $1$-dimensional. Applying the rank-nullity theorem, the column span of $\mathbf{Q}-\mathbf{I}$ is $(m-1)$-dimensional. Since $(\mathbf{Q}-\mathbf{I})\bm{1}_{m}=\mathbf{Q}\bm{1}_{m}-\mathbf{I}\bm{1}_{m}=\bm{1}_{m}-\bm{1}_{m}=\bm{0}_{m}$, each column of $\mathbf{Q}-\mathbf{I}$ is redundant, i.e., removing any column of $\mathbf{Q}-\mathbf{I}$ does not change the column span.
Next, observe that $\bm{\pi}\bm{1}_{m}=1\neq 0,$ implying that $\bm{\pi}$ is not orthogonal to $\bm{1}_{m}$. In contrast, each vector $\mathbf{v}$ in the column span of $\mathbf{Q}-\mathbf{I}$ is of the form $\mathbf{v}=(\mathbf{Q}-\mathbf{I})\mathbf{x}$ for some $\mathbf{x}\in\mathcal{R}^{m}$. Thus $\bm{\pi}\mathbf{v}=\bm{\pi}(\mathbf{Q}-\mathbf{I})\mathbf{x}=(\bm{\pi}\mathbf{Q}-\bm{\pi}\mathbf{I})\mathbf{x}=(\bm{\pi}-\bm{\pi})\mathbf{x}=\bm{0}_{m}^{T}\mathbf{x}=0$, implying that $\bm{\pi}$ is orthogonal to $\mathbf{v}$. Therefore, $\bm{1}_{m}$ is not in the column span of $\mathbf{Q}-\mathbf{I}$.
Combining these two observations, replacing the first column of $\mathbf{Q}-\mathbf{I}$ with $\bm{1}_{m}$ increases the rank of $\mathbf{Q}-\mathbf{I}$ by exactly one. Hence, $\mathbf{M}$ has rank $m$. This shows invertibility, and the proof of (a) follows.
To prove (b) and (c) observe that
$\bm{\pi}\begin{bmatrix}y^{int}\\ y^{int}\\ \vdots\\ y^{int}\\ \end{bmatrix}=\bm{\pi}\bm{1}_{m}y^{int}=(1)y^{int}=y^{int},$
and that for all $\mathbf{x}\in\mathcal{R}^{m-1}$,
$\bm{\pi}\begin{bmatrix}f_{0}(\mathbf{x},S_{0})\\ f_{1}(\mathbf{x},S_{1})\\ \vdots\\ f_{m-1}(\mathbf{x},S_{m-1})\\ \end{bmatrix}$ $=\bm{\pi}\begin{bmatrix}\ell(S_{0})\\ \ell(S_{1})\\ \vdots\\ \ell(S_{m-1})\\ \end{bmatrix}+\bm{\pi}(\mathbf{Q}-\mathbf{I})\begin{bmatrix}0\\ x_{1}\\ \vdots\\ x_{m-1}\\ \end{bmatrix},$
$=\begin{bmatrix}\pi_{0}(\mathbf{S})\\ \pi_{1}(\mathbf{S})\\ \vdots\\ \pi_{m-1}(\mathbf{S})\\ \end{bmatrix}^{T}\begin{bmatrix}\ell(S_{0})\\ \ell(S_{1})\\ \vdots\\ \ell(S_{m-1})\\ \end{bmatrix}+{\bf 0}_{m}\begin{bmatrix}0\\ x_{1}\\ \vdots\\ x_{m-1}\\ \end{bmatrix},$
$=\sum_{k=0}^{m-1}\ell(S_{k})\cdot\pi_{k}(\mathbf{S}),$
$={\rm cost}(\mathbf{S}).$
Applying these observations by setting $\mathbf{x}=\mathbf{x}^{int}$ and left-multiplying by $\bm{\pi},$
$y^{int}=\bm{\pi}\begin{bmatrix}y^{int}\\ y^{int}\\ \vdots\\ y^{int}\\ \end{bmatrix}=\bm{\pi}\begin{bmatrix}f_{0}(\mathbf{x}^{int},S_{0})\\ f_{1}(\mathbf{x}^{int},S_{1})\\ \vdots\\ f_{m-1}(\mathbf{x}^{int},S_{m-1})\\ \end{bmatrix}={\rm cost}(\mathbf{S}).$
Applying these observations again, it follows that
$\forall\mathbf{x}\in\mathcal{R}^{m-1},\quad y^{int}=cost(\mathbf{S})=\bm{\pi}\begin{bmatrix}f_{0}(\mathbf{x},S_{0})\\ f_{1}(\mathbf{x},S_{1})\\ \vdots\\ f_{m-1}(\mathbf{x},S_{m-1})\\ \end{bmatrix}=\sum_{k=0}^{m-1}f_{k}(\mathbf{x},S_{k})\cdot\pi_{k}(\mathbf{S})$
proving (b). Since the transition probabilities $\pi(S_{k})\geq 0$ are non-negative,
$\forall\mathbf{x}\in\mathcal{R}^{m-1},\quad y^{int}\geq\min_{k\in[m]}f_{k}(\mathbf{x},S_{k})\cdot\sum_{k=0}^{m-1}\pi(S_{k})=\min_{k\in[m]}f_{k}(\mathbf{x},S_{k}),$
proving (c).
(d) follows by observing that
$\forall\ \mathbf{x}\in\mathcal{R}^{m-1},\quad y^{int}=\min_{k\in[m]}f_{k}(\mathbf{x},S_{k})\geq\min_{k\in[m]}g_{k}(\mathbf{x})=h(\mathbf{x}).$
5.2 Proof of Lemma 3.7
Proof.
To prove (a) set $\mathbf{S}_{|P}(\hat{\mathbf{x}})=\left(S_{0|P}(\hat{\mathbf{x}}),S_{1|P}(\hat{\mathbf{x}}),\dots,S_{m-1|P}(\hat{\mathbf{x}})\right).$ Assume that, $\forall k\in P,$ $g_{k|P}(\hat{\mathbf{x}})=f_{k}(\hat{\mathbf{x}},S_{k|P}(\hat{\mathbf{x}}))=\hat{y}.$
By definition, $S_{k|P}(\hat{\mathbf{x}})\in{\mathbb{S}}_{k|P}$ cannot transition to any $S_{j}\in{\mathbb{S}}_{j}$ where $j\not\in P$. This implies that
$\forall k\not\in P,\ \quad\pi_{k}(\mathbf{S}_{|P}(\hat{\mathbf{x}}))=0\quad\implies\quad\sum_{k\in P}\pi_{k}(\mathbf{S}_{|P}(\hat{\mathbf{x}}))=1.$
Lemma 3.1 (b) then implies
${\rm cost}(\mathbf{S}_{|P}(\hat{\mathbf{x}}))$ $=$ $\sum_{k=0}^{m-1}f_{k}(\hat{\mathbf{x}},S_{k|P}(\hat{\mathbf{x}}))\cdot\pi_{k}(\mathbf{S}_{|P}(\hat{\mathbf{x}}))$
$=$ $\sum_{k\in P}f_{k}(\hat{\mathbf{x}},S_{k|P}(\hat{\mathbf{x}}))\cdot\pi_{k}(\mathbf{S}_{|P}(\hat{\mathbf{x}}))$
$=$ $\sum_{k\in P}\hat{y}\cdot\pi_{k}(\mathbf{S}_{|P}(\hat{\mathbf{x}}))$
$=$ $\hat{y}.$
Thus, $\mathbf{S}_{|P}(\hat{\mathbf{x}})$ is a minimum-cost Markov chain.
To prove (b) first note that, we are given that $\forall k,$ $q_{0}(S^{*}_{k})>0.$ Thus, $\pi(S^{*}_{0})>0$ and $\{0\}\subseteq P^{*}$, as required.
The Markov chain starts in state $S^{*}_{0}$ where $0\in P^{*}$. We claim that if $k\in P^{*}$ and $j\not\in P^{*}$ then $q_{j}(S^{*}_{k})=0.$ If not, $\pi_{j}(\mathbf{S}^{*})\geq\pi_{k}(\mathbf{S}^{*})\cdot q_{j}(S^{*}_{k})>0$, contradicting that $\pi_{j}(\mathbf{S}^{*})=0.$ Thus, $S^{*}_{k}\in{\mathbb{S}}_{k|P^{*}}$ for all $k\in P^{*}$.
Now, suppose that $P^{*}\subseteq P.$ Then, for all $k\in P^{*}$,
$f_{k}(\hat{\mathbf{x}},S^{*}_{k})\geq g_{k|P^{*}}(\hat{\mathbf{x}})\geq g_{k|P}(\hat{\mathbf{x}})\geq g_{k}(\hat{\mathbf{x}})\geq h(\hat{\mathbf{x}})=\hat{y}.$ (9)
On the other hand, using Lemma 3.1 (b) and the fact that $\mathbf{S}^{*}$ is a minimum-cost Markov chain,
$\hat{y}={\rm cost}(\mathbf{S}^{*})=\sum_{k\in P^{*}}f_{k}(\hat{\mathbf{x}},S^{*}_{k})\cdot\pi_{k}(\mathbf{S}^{*})\geq\hat{y}\sum_{k\in P^{*}}\pi_{k}(\mathbf{S}^{*})=\hat{y}.$ (10)
The left and right hand sides of (10) are the same and $\forall k\in P^{*},\,\pi_{k}(\mathbf{S}^{*})>0,$ so $\forall k\in P^{*},\,f_{k}(\hat{\mathbf{x}},S^{*}_{k})=\hat{y}$.
This in turn forces all the inequalities in (9) to be equalities, i.e., for all $k\in P^{*}$, $g_{k|P}(\hat{\mathbf{x}})=\hat{y}.$
Thus, $P^{*}\subseteq\{k\in P:g_{k|P}(\hat{\mathbf{x}})=\hat{y}\}$, proving (b). ∎
6 A Polynomial Time Algorithm for binary AIFV-$m$ Coding.
This second part of the paper introduces binary AIFV-$m$ codes and then applies Lemma 4.6 to find a minimum-cost code of this type in time polynomial in $n$, the number of source words to be encoded, and $b,$ the number of bits required to state the probability of any source word.
The remainder of this section defines binary AIFV-$m$ codes and describes how they are a special case of the minimum-cost Markov chain problem.
Section 7 describes how to apply Lemma 4.6.
This first requires showing that $\varphi$ is polynomial in $n$ and $b$, which will be straightforward. It also requires identifying a hyperrectangle $\bf R$ that contains a highest point $(\mathbf{x}^{*},y^{*})\in\mathbb{H}$ and for which $t^{\prime}_{{\mathbb{S}}}({\bf R})$ is polynomial in $n$. That is, $\forall\mathbf{x}\in{\bf R},\forall P\in\mathcal{P},$ $\mathbf{S}_{|P}(\mathbf{x})$ can be calculated in polynomial time.
As mentioned at the start of Section 4, as part of improving the running time of the iterative algorithm, [16, 9, 22, 11] showed that, for ${\bf R}=[0,1]^{m-1},$ $t_{{\mathbb{S}}}({\bf R})$ is polynomial time. As will be discussed in Section 7.2, the algorithms there can be easily modified to show that $t^{\prime}_{{\mathbb{S}}}({\bf R})=O\left(t_{{\mathbb{S}}}({\bf R})\right)$.
Corollary 3.4 only tells us that there exists some $\mathbf{x}^{*}$ such that $(\mathbf{x}^{*},h(\mathbf{x}^{*}))$ is a highest point in $\mathbb{H}$. In order to use Lemma 4.6, we will need to show that there exists such an $\mathbf{x}^{*}\in[0,1]^{m-1}$. Proving this is the most cumbersome part of the proof. It combines a case-analysis of the tree structures of AIFV-$m$ trees with the Poincare-Miranda theorem to show that the functions $g_{k}(\mathbf{x}),$ $k\in[m]$ must all mutually intersect at some point $\mathbf{x}^{*}\in[0,1]^{m-1}.$ From Lemma 3.3, $\left(\mathbf{x}^{*},h\left(\mathbf{x}^{*}\right)\right)\in\mathbb{H}$ and is therefore the optimum point needed. Section 8 develops the tools required for this analysis.
6.1 Background
Consider a stationary memoryless source with alphabet $\Sigma=\{\sigma_{1},\sigma_{2},\dots,\sigma_{n}\}$ in which symbol $\sigma_{i}$ is generated with probability $p_{i}$. Let ${\mathcal{M}}=\alpha_{1}\,\alpha_{2}\,\alpha_{3}\,\ldots$ be a message generated by the source.
Binary compression codes encode each $\sigma_{i}$ in ${\mathcal{M}}$ using a binary codeword. Huffman codes are known to be “optimal” such codes. More specifically, they are Minimum Average-Cost Binary Fixed-to-Variable Instantaneous codes. “Fixed-to-Variable” denotes that the binary codewords corresponding to the different $\sigma_{i}$ can have different lengths. “Instantaneous”, that, in a bit-by-bit decoding process, the end of a codeword is recognized immediately after its last bit is scanned. The redundancy of a code is the difference between its average-cost and the Shannon entropy $-\sum_{i}p_{i}\log_{2}p_{i}$ of the source. Huffman codes can have worst case redundancy of $1$
Huffman codes are often represented by a coding tree, with the codewords being the leaves of the tree. A series of recent work [4, 5, 6, 14, 16, 22, 23] introduced Binary Almost-Instantaneous Fixed-to-Variable-$m$ (AIFV-$m$) codes. Section 6.2 provides a complete definition as well as examples. These differ from Huffman codes in that they use $m$ different coding trees. Furthermore, decoding might require reading ahead $m$ bits before knowing that the end of a codeword has already been reached (hence “almost”-instantaneous). Since AIFV-$m$ codes include Huffman coding as a special case, they are never worse than Huffman codes. Their advantage is that, at least for $m\leq 5,$ they have worst-case redundancy $1/m$ [14, 6], beating Huffman coding. [^6] Huffman coding with blocks of size $m$ will also provide worst case redundancy of $1/m.$ But the block source coding alphabet, and thus the Huffman code dictionary, would then have size $\Theta(n^{m}).$ In contrast, AIFV-$m$ codes have dictionary size $\Theta(mn).$
Historically, AIFV-$m$ codes were preceded by $K$-ary Almost-Instantaneous FV ($K$-AIFV) codes, introduced in [21]. $K$-ary AIFV codes used a $K$ character encoding alphabet; for $K>2$, the procedure used $K-1$ coding trees and had a coding delay of $1$ bit. For $K=2,$ it used 2 trees and had a coding delay of $2$ bits.
Binary AIFV-$m$ codes were introduced later in [14]. These are binary codes that are comprised of an $m$-tuple of binary coding trees and have decoding delay of at most $m$ bits. The binary AIFV-$2$ codes of [14] are identical to the $2$-ary AIFV codes of [21].
Constructing optimal777 An “optimal” $K$-AIFV or AIFV-$m$ code is one with minimum average encoding cost over all such codes. This will be formally specified later in Definition 6.8. $K$-AIFV or binary AIFV-$m$ codes is much more difficult than constructing Huffman codes. [23] described an iterative algorithm for constructing optimal binary AIFV-$2$ codes. [22] generalized this and proved that, for $m=2,3,4,5$, under some general assumptions, this algorithm would terminate and, at termination would produce an optimal binary AIFV-$m$ code. The same was later proven for $m>5$ by [4] and [22]. This algorithm was later generalized to solve the Minimun Cost Markov chain problem in [5] and is the iterative algoithm referenced in Section 2.3.
6.2 Code Definitions, Encoding and Decoding
Note: We assume $\sum_{i}p_{i}=1$ and that each $p_{i}$ can be represented using $b$ bits, i.e., each probability is an integral multiple of $2^{-b}$. The running time of our algorithm will, for fixed $m,$ be polynomial in $b$ and $n$, i.e., weakly polynomial.
A binary AIFV-$m$ code will be an $m$-tuple $\mathbf{T}=(T_{0},T_{1},\dots,T_{m-1})$ of $m$ binary code trees satisfying Definitions 6.1 and 6.2 below. Each $T_{i}$ contains $n$ codewords. Unlike in Huffman codes, codewords can be internal nodes.
Definition 6.1 (Node Types in a Binary AIFV-$m$ Code [14]).
Figure 4. Edges in an AIFV-$m$ code tree are labelled as 00-edges or $1$-edges. If node $v$ is connected to its child node $u$ via a 00-edge ($1$-edge) then $u$ is $v$’s 00-child ($1$-child). We will often identify a node interchangeably with its associated (code)word. For example, $100^{2}$ 10 is the node reached by following the edges $0,0,1,0$ down from the root. Following [14], the nodes in AIFV-$m$ code trees can be classified as being exactly one of 3 different types:
  • Complete Nodes. A complete node has two children: a 00-child and a $1$-child. A complete node has no source symbol assigned to it.
  • Intermediate Nodes. An intermediate node has no source symbol assigned to it and has exactly one child. An intermediate node with a 00-child is called an intermediate-00 node; with a $1$-child is called an intermediate-$1$ node.
  • Master Nodes. A master node has an assigned source symbol and at most one child node. Master nodes have associated degrees:
    • a master node of degree $k=0$ is a leaf.
[…]
Figure 4: Node types in a binary AIFV-$3$ code tree: complete node ($C$), intermediate-00 and intermediate-$1$ nodes ($I_{0}$, $I_{1}$), master nodes of degrees $0,1,2$ ($M_{0}$, $M_{1}$, $M_{2}$), ($M_{0}$ is a leaf)and non-intermediate-00 nodes (${N_{0}}$). The ${N_{0}}$ nodes can be complete, master, or intermediate-$1$ nodes, depending upon their location.
Binary AIFV-$m$ codes are now defined as follows:
Definition 6.2 (Binary AIFV-$m$ Codes [14]).
See Figure 5. Let $m\geq 2$ be a positive integer. A binary AIFV-$m$ code is an ordered $m$-tuple of $m$ code trees $(T_{0},T_{1},\dots,T_{m-1})$ satisfying the following conditions:
  1. 1.
    Every node in each code tree is either a complete node, an intermediate node, or a master node of degree $k$ where $0\leq k<m$.
  2. 2.
    For $k\geq 1$, the code tree $T_{k}$ has an intermediate-$1$ node connected to the root by exactly $k$ 00-edges, i.e., the node $0^{k}$ is an intermediate-$1$ node.
Consequences of the Definitions:
  1. (a)
    Every leaf is a master node of degree 00. In particular, this implies that every code tree contains at least one master node of degree $0$
  2. (b)
    Definition 6.2, and in particular Condition (2), result in unique decodability (proven in [14]).
  3. (c)
    For $k\not=1,$ the root of a $T_{k}$ tree is permitted to be a master node. If a root is a master node, the associated codeword is the empty string (Figure 5)! The root of a $T_{1}$ tree cannot be a master node.
  4. (d)
    The root of a $T_{k}$ tree may be an intermediate-00 node.
  5. (e)
    For $k>0,$ every $T_{k}$ tree must contain at least one intermediate-$1$ node, the node $0^{k}.$ A $T_{0}$ tree might not contain any intermediate-$1$ node.
    For $k>0,$ the root of a $T_{k}$ tree cannot be a intermediate-$1$ node. The root of a $T_{0}$ tree is permitted to be an intermediate-$1$ node (but see Lemma 8.7).
Figure 5: Example binary AIFV-$3$ code for source alphabet $\left\{a,b,c,d\right\}.$ The small filled nodes are complete nodes; the small striped nodes are intermediate-$1$ nodes and the small empty nodes are intermediate-00 ones. The large nodes are master nodes with their assigned source symbols. They are labelled to their sides as $M_{i}$ nodes, indicating that they are master-$i$ nodes. Note that $T_{2}$ encodes $a$, which is at its root, with an empty string!
We now describe the encoding and decoding procedures. These are illustrated in Figures 6 and 7.
Procedure 6.3 (Encoding of a Binary AIFV-$m$ Code).
A source sequence ${\mathcal{M}}=\alpha_{1}\alpha_{2}\dots$ is encoded as follows: Set $T=T_{0}$ and $i=1.$
1. Encode $\alpha_{i}$ using $T$
2. Let $k$ be the index such that $\alpha_{i}$ is encoded using a degree-$k$ master node in $T$
3. Set $T=T_{k};$ $i=i+1$
4. Goto line 1
[…]
Figure 6: Encoding $cbab$ using the binary AIFV-$3$ code in Figure 5. The first $c$ is encoded via $T_{0}$ as “$000$” using a degree-00 master node (leaf). Thus the first $b$ is encoded via $T_{0}$ as “$1$” using a degree-$2$ master node. Thus the first $a$ is encoded via $T_{2}$ as the empty string ($\epsilon$) using a degree-$1$ master node. Thus the second $b$ is encoded via $T_{1}$ as “$010$” using a degree-00 master node. $\ldots$.… .
Procedure 6.4 (Decoding of a Binary AIFV-$m$ Code).
Let $\beta$ be a binary string that is the encoded message. Set $T=T_{0}$ and $i=1.$
1. Let $\beta_{i}$ be longest prefix of $\beta$ that corresponds to a path from the root of $T$
to some master node $M$ in $T$
2. Let $k$ be the degree of $M$ (as a master node) in $T$
3. Set $\alpha_{i}$ to be the source symbol assigned to $\beta_{i}$ in $T$
4. Remove $\beta_{i}$ from the start of $\beta.$
5. Set $T=T_{k};$ $i=i+1$
6. Goto line 1
Note that in order to identify $\beta_{i}$, line 1 might require reading a few bits after the end of $\beta_{i}$. The number of extra bits that must be read is known as the delay.
[…]
Figure 7: Decoding $0001010$ using the binary AIFV-$3$ code in Figure 5.
Theorem 6.5 ([14], Theorem 3).
Binary AIFV-$m$ codes are uniquely decodable with delay at most $m$.
6.3 The cost of AIFV-$m$ codes
$\begin{array}[]{||cc||ccc|ccc|ccc||}\hline\cr i&\sigma_{i}&\ell(T_{0},\sigma_{i})&d(T_{0},\sigma_{i})&w(T_{0},\sigma_{1})&\ell(T_{1},\sigma_{i})&d(T_{1},\sigma_{i})&w(T_{1},\sigma_{1})&\ell(T_{2},\sigma_{i})&d(T_{2},\sigma\_i)&w(T_{2},\sigma_{1})\\ \hline\cr 1&a&1&1&0&1&2&1&0&1&\epsilon\\ \hline\cr 2&b&1&2&1&3&0&010&4&0&0010\\ \hline\cr 3&c&3&0&000&3&0&011&5&0&00110\\ \hline\cr 4&d&4&0&1000&4&0&1000&5&0&00111\\ \hline\cr\end{array}$
$\begin{array}[]{|c|ccc|}\hline\cr{j}&\mathcal{M}_{j}(T_{0})&\mathcal{M}_{j}(T_{1})&\mathcal{M}_{j}(T_{2})\\ \hline\cr 0&\{3,4\}&\{2,3,4\}&\{2,3,4\}\\ 1&\{1\}&\emptyset&\{1\}\\ 2&\{2\}&\{1\}&\emptyset\\ \hline\cr\end{array}\quad\begin{array}[]{|c|ccc|}\hline\cr{j}&q_{j}(T_{0})&q_{j}(T_{1})&q_{j}(T_{2})\\ \hline\cr 0&p_{3}+p_{4}&p_{2}+p_{3}+p_{4}&p_{2}+p_{3}+p_{4}\\ 1&p_{1}&0&p_{1}\\ 2&p_{2}&p_{1}&0\\ \hline\cr\end{array}$
Figure 8: Example of the values introduced in Definition 6.7 when calculated for the AIFV-$3$ code given in Figure 5. $w(T_{k},\sigma_{i})$ in the first table denotes the codeword for $\sigma_{i}$ in code tree $T_{k}.$ Note that the rightmost table on the second row is the transition matrix $\mathbf{Q}(\mathbf{T})$ for the code.
Definition 6.6.
Let $\mathcal{T}_{k}(m,n)$ denote the set of all possible type-$k$ trees that can appear in a binary AIFV-$m$ code on $n$ source symbols. $T_{k}$ will be used to denote a tree $T_{k}\in\mathcal{T}_{k}(m,n)$. Set $\mathcal{T}(m,n)=\bigtimes_{k=0}^{m-1}\mathcal{T}_{k}(m,n)$.
$\mathbf{T}=(T_{0},\ldots,T_{m-1})\in\mathcal{T}(m,n)$ will be called a binary AIFV-$m$ code.
Definition 6.7.
Figure 8. Let $T_{k}\in\mathcal{T}_{k}(m,n)$ and $\sigma_{i}$ be a source symbol.
  • $\ell(T_{k},\sigma_{i})$ denotes the length of the codeword in $T_{k}$ for $\sigma_{i}$.
  • $d(T_{k},\sigma_{i})$ denotes the degree of the master node in $T_{k}$ assigned to $\sigma_{i}$.
  • $\ell(T_{k})$ denotes the average length of a codeword in $T_{k}$, i.e.,
    $\ell(T_{k})=\sum_{i=1}^{n}\ell(T_{k},\sigma_{i})\cdot p_{i}.$
  • $\mathcal{M}_{j}(T_{k})=\{i\in\{1,2,\ldots,n\}:\,d(T_{k},\sigma_{i})=j\}$ is the set of indices of source nodes that are assigned master nodes of degree $j$ in $T_{k}.$ Set
    ${\bf q}\left(T_{k})=(q_{0}(T_{k}),\ldots,q_{m-1}(T_{k})\right)\quad\mbox{ where }\quad\forall j\in[m],\,\ q_{j}(T_{k})=\sum_{i\in\mathcal{M}_{j}(T_{k})}p_{i}.$
    $\sum_{j\in m}q_{j}(T_{k})=1,$ so ${\bf q}(T_{k})$ is a probability distribution.
If a source symbol is encoded using a degree-$j$ master node in $T_{k}$, then the next source symbol will be encoded using code tree $T_{j}$. Since the source is memoryless, the transition probability of encoding using code tree $T_{j}$ immediately after encoding using code tree $T_{k}$ is $q_{j}(T_{k})$.
Figure 9: Markov chain corresponding to AIFV-$3$ code in Figure 5. Note that $T_{1}$ contains no degree 1 master node, so there is no edge from $T_{1}$ to $T_{1}.$ Similarly, $T_{2}$ contains no degree 2 master node, so there is no edge from $T_{2}$ to $T_{2}.$
This permits viewing the process as a Markov chain whose states are the code trees. Figure 9 illustrates an example.
From Consequence (a) following Definition 6.2, $\forall k\in[m],$ every $T_{k}\in\mathcal{T}_{k}(m,n)$ contains at least one leaf, so $q_{0}(T_{k})>0$. Thus, as described in Section 2.1, this implies that the associated Markov chain is a unichain whose unique recurrence class contains $T_{0}$ and whose associated transition matrix $\mathbf{Q}$ has a unique stationary distribution $\bm{\pi}$. Thus the Markov chain associated with any $\mathbf{T}\in\mathcal{T}(m,n)$ is permissible.
Definition 6.8.
Let $\mathbf{T}=(T_{0},\ldots,T_{m-1})\in\mathcal{T}(m,n)$ be some AIFV-$m$ code, $\mathbf{Q}(\mathbf{T})$ be the transition matrix of the associated Markov chain and
$\bm{\pi}(\mathbf{T})=(\pi_{0}(\mathbf{T}),\ldots,\pi_{m-1}(\mathbf{T}))$
be $\mathbf{Q}(\mathbf{T})$’s associated unique stationary distribution. Then the average cost of the code is the average length of an encoded symbol in the limit, i.e.,
${\rm cost}(\mathbf{T})=\sum_{k=0}^{m-1}\ell(T_{k})\cdot\pi_{k}(\mathbf{T}).$
Definition 6.9 (The Binary AIFV-$m$ Code problem).
Construct a binary AIFV-$m$ code $\mathbf{T}\in\mathcal{T}(m,n)$ with minimum ${\rm cost}(\mathbf{T}),$ i.e.,
${\rm cost}(\mathbf{T})=\min_{\mathbf{T}^{\prime}\in\mathcal{T}(m,n)}{\rm cost}(\mathbf{T}^{\prime}).$
This problem is exactly the minimum-cost Markov Chain problem introduced in Section 2.1 with ${\mathbb{S}}_{k}=\mathcal{T}_{k}(m,n)$. As discussed in the introduction to Section 4, this was originally solved in exponential time by using an iterative algorithm.
7 Using Lemma 4.6 to derive a polynomial time algorithm for binary AIFV-$m$ coding
Because the minimum-cost binary AIFV-$m$ coding problem is a special case of the minimum-cost Markov chain problem, Lemma 4.6 can be applied to derive a polynomial time algorithm. In the discussion below working through this application, $T_{k}$ will denote, interchangeably, both a type-$k$ tree and a type-$k$ state with transition probabilities $q_{j}(T_{k})$ and cost $\ell(T_{k}).$ For example, when writing $f_{k}(\mathbf{x},T_{k})$ (as in Definition 2.4), $T_{k}$ will denote the corresponding Markov chain state and not the tree.
Applying Lemma 4.6 requires showing that, for fixed $m,$ the $\varphi$ and $t^{\prime}_{{\mathbb{S}}}({\bf R})$ parameters in its statement are polynomial in $b$ and $n$.
7.1 Showing that $\varphi$ is polynomial in $b,n$ for AIFV-$m$ coding
Recall that $\varphi$ is the maximum number of bits needed to represent the coefficients of any linear inequality defining a constraint of $\mathbb{H}.$
Showing that $\varphi$ is polynomial in $b$ and $n$ is not difficult but will require the following fact proven later in Corollary 8.9:
            For all $T_{k}\in\mathcal{T}_{k}(m,n),$ the height of $T_{k}$ is at most $n(m+1).$
Lemma 7.1.
${\mathbb{H}}$ is defined by inequalities of size $O(mb+\log n+\log m)$ where $b$ is the maximum number of bits needed to encode any of the $p_{i}.$
Proof.
Note that the definition of ${\mathbb{H}}$ can be equivalently written as
${\mathbb{H}}$ $=$ $\left\{(\mathbf{x},y)\,|\,(\mathbf{x}\in\mathcal{R}^{m-1},\,y\in\mathcal{R},\,y\leq h(\mathbf{x})\right\}$
$=$ $\bigcap_{k\in[m]}\ \bigcap_{T_{k}\in\mathcal{T}_{k}(m,n)}\left\{(\mathbf{x},y)\,|\,\mathbf{x}\in\mathcal{R}^{m-1},\,y\in\mathcal{R},\,y\leq f_{k}(\mathbf{x},T_{k})\right\}$
$=$ $\bigcap_{k\in[m]}\ \bigcap_{T_{k}\in\mathcal{T}_{k}(m,n)}\left\{(\mathbf{x},y)\,|\,\mathbf{x}\in\mathcal{R}^{m-1},\,y\in\mathcal{R},\,y+x_{k}-\sum_{j=1}^{m-1}q_{j}(T_{k})\cdot x_{j}\leq\ell(T_{k})\right\}$
where we set $x_{0}\equiv 0$ to provide notational consistency between $k=0$ and $k>0.$ Thus, the linear inequalities defining ${\mathbb{H}}$ are of the form
$y+x_{k}-\sum_{j=1}^{m-1}q_{j}(T_{k})\cdot x_{j}\leq\ell(T_{k}).$ (11)
Since each $p_{i}$ can be represented with $b$ bits, $p_{i}=P_{i}2^{-b}$ for some integral $P_{i}\in(0,2^{b}).$ This implies that $q_{j}(T_{k})=Q_{j,k}2^{-b}$ for some integral $Q_{j,k}\in[0,2^{b}].$ So the size of each $q_{j}(T_{k})$ is $\leq 2b.$
Corollary 8.9, to be proven later, states that the height of $T_{k}$ is bounded by $n(m+1).$ Thus each $\ell(T_{k},\sigma_{i})\leq n(m+1)$ for all $k,i$ and $\ell(T_{k})=\sum_{i=1}^{n}\ell(T_{k},\sigma_{i})\cdot p_{i}$ can be written as $L_{k}2^{-b}$ for some integral $L_{k}\in\left[1,(n(m+1))2^{b}\right].$ Thus, $\ell(T_{k})$ has size $O(b+\log n+\log m).$ Thus, the size of every inequality (11) is at most $O(mb+\log n+\log m).$
Recall that $m$ is considered fixed. Thus $\varphi=O(b+\log n).$
7.2 Finding Appropriate $\bf R$ and Showing that $t_{{\mathbb{S}}}({\bf R})$ and $t^{\prime}_{{\mathbb{S}}}({\bf R})$ are Polynomial in $b,n$ for AIFV-$m$ Coding
Recall that ${\mathbb{S}}=\mathcal{T}(m,n).$ Fix $m$ and ${\bf R}=[0,1]^{m-1}.$
As discussed at the starts of Section 4 and 6, there are dynamic programming algorithms that, for $m=2,$ give $t_{{\mathbb{S}}}({\bf R})=O(n^{3})$ [9] and for $m>2,$  $t_{{\mathbb{S}}}({\bf R})=O(n^{m+2})$ [11]. For $\mathbf{x}\not\in{\bf R},$ the best known algorithms for calculating $\mathbf{S}(\mathbf{x})$ use integer linear programming and run in exponential time.
In deriving the polynomial time binary search algorithm for $m=2$, [10] proved that $\mathbf{x}^{*}\in[0,1],$ and could therefore use the $O(n^{3})$ DP time algorithm for $\mathbf{S}(\mathbf{x})$ as a subroutine. We need to prove something similar for $m>2.$
The main tool used will be the following highly technical lemma whose proof is deferred to the next section.
Lemma 7.2.
Let $m$ be fixed, $n\geq 2^{m-1},$ $\mathbf{x}\in{[0,1]^{m-1}}$ and $k\in\{1\,\ldots,\,m-1\}.$ Then
  • If $x_{k}=0,$ $g_{0}(\mathbf{x})-g_{k}(\mathbf{x})\leq 0.$
  • If $x_{k}=1,$ $g_{k}(\mathbf{x})-g_{0}(\mathbf{x})\leq 0.$
The proof also needs a generalization of the intermediate-value theorem:
Theorem 7.3 (Poincaré-Miranda Theorem [19]).
Let $f_{1},f_{2},\dots,f_{r}:[0,1]^{r}\to\mathcal{R}$ be $r$ continuous functions of $r$ variables $x_{1},x_{2},\dots,x_{r}\in[0,1]$ such that for all indices $i$, $x_{i}=0$ implies that $f_{i}\leq 0$ and $x_{i}=1$ implies that $f_{i}\geq 0$. It follows that there exists a point $(x_{1},x_{2},\dots,x_{r})\in[0,1]^{r}$ such that $\forall i,\,f_{i}(x_{1},\ldots,x_{r})=0$.
Combining the two yields:
Lemma 7.4.
Let $m$ be fixed and $n\geq 2^{m-1}.$ Then there exists $\mathbf{x}^{*}\in[0,1]^{m-1}$ satisfying
$g_{0}(\mathbf{x}^{*})=g_{1}(\mathbf{x}^{*})=\dots=g_{m-1}(\mathbf{x}^{*}).$ (12)
Proof.
Set $r\leftarrow m-1$ and $f_{k}\leftarrow g_{0}-g_{k}$ for all $k\in\{1,\dots,m-1\}$. From Lemma 7.2, if $x_{k}=0$ then $f_{k}(\mathbf{x})\leq 0$ and if $x_{k}=1$ then $f_{k}(\mathbf{x})\geq 0.$ The Poincaré-Miranda theorem then immediately implies the existence of $\mathbf{x}^{*}\in[0,1]^{m-1}$ such that, $\forall k\in\{1,\dots,m-1\},\,g_{0}(\mathbf{x}^{*})=g_{k}(\mathbf{x}^{*})$, i.e., (12). ∎
Lemma 7.4 combined with Lemma 3.3, immediately show that if $n\geq 2^{m-1},$ there exists $\mathbf{x}^{*}\in[0,1]^{m-1}$ satisfying $h(\mathbf{x}^{*})=\max\{y^{\prime}\,:\,(\mathbf{x}^{\prime},y^{\prime})\in\mathbb{H}\}$.
As noted earlier, for $\mathbf{x}\in{\bf R}=[0,1]^{m-1},$ there are dynamic programming algorithms for calculating $\mathbf{S}(\mathbf{x})$ in $O(n^{3})$ time when $m=2$ [9] and $O(n^{m+2})$ when $m>2$ [11]. Thus $t_{{\mathbb{S}}}({\bf R})=O(n^{3})$ for $m=2$ and $t_{{\mathbb{S}}}({\bf R})=O(n^{m+2})$ for $m>2.$
Those dynamic programming algorithms work by building the $T_{k}$ trees top-down. The status of nodes on the bottom level of the partially built tree, i.e., whether they are complete, intermediate nodes or master nodes of a particular degree, is left undetermined. One step of the dynamic programming algorithm then determines (guesses) the status of those bottom nodes and creates a new bottom level of undetermined nodes. It is easy to modify this procedure so that nodes are only assigned a status within some given $P\in\mathcal{P}.$ The modified algorithms would then calculate $\mathbf{S}_{|P}(\mathbf{x})$ in the same running time as the original algorithms, i.e., $O(n^{3})$ time when $m=2$ and $O(n^{m+2})$ when $m>2$. Thus, for $\mathbf{x}\in{\bf R}$, $t^{\prime}_{{\mathbb{S}}}({\bf R})=O\left(t_{{\mathbb{S}}}({\bf R})\right).$
7.3 The Final Polynomial Time Algorithm
Fix $m$ and set ${\bf R}=[0,1]^{m-1}$. Since $m$ is fixed, we may assume that $n\geq 2^{m-1}.$ For smaller $n,$ the problem can be solved in $O(1)$ time by brute force.
In the notation of Lemma 4.6, Section 7.1 shows that $\varphi=O(b+\log n)$ where $b$ is the maximum number of bits needed to encode any of the $p_{i}.$ Section 7.2 shows that $t^{\prime}_{{\mathbb{S}}}({\bf R})=O(n^{3})$ when $m=2$ and $O(n^{m+2})$ when $m>2$ and there always exists $\mathbf{x}^{*}\in{\bf R}$ satisfying $h(\mathbf{x}^{*})=\max\{y^{\prime}\,:\,(\mathbf{x}^{\prime},y^{\prime})\in\mathbb{H}\}$.
Then, from Lemma 4.6, the binary AIFV-$m$ coding problem can be solved in time polynomially bounded by $\varphi$ and $t^{\prime}_{S}({\bf R}),$ i.e., weakly polynomial in the input.
8 Proof of Lemma 7.2 and Corollary 8.9
The polynomial running time of the algorithm rested upon the correctness of the technical Lemma 7.2 and Corollary 8.9. The proof of Lemma 7.2 requires deriving further technical properties of AIFV-$m$ trees. Corollary 8.9 will be a consequence of some of these derivations.
The main steps of the proof of Lemma 7.2 are:
  • generalize binary AIFV-$m$ codes trees to extended binary code trees;
  • prove Lemma 7.2 for these extended trees;
  • convert this back to a proof of the original lemma.
We first introduce the concept of extended binary code trees.
Definition 8.1 (Extended Binary AIFV-$m$ Codes).
Fix $k\in[m].$ An extended binary AIFV-$m$ code tree $T^{ex}_{k}$ is defined exactly the same as a $T_{k}$ except that it is permitted to have an arbitrary number of leaves, i.e., master nodes of degree $0,$ assigned the empty symbol $\epsilon.$ The $\epsilon$-labelled leaves are given source probabilities $0$ [^8] Since $T^{ex}_{k}$ satisfies Definitions 6.1 and 6.2, just for a larger number of master nodes, it also satisfies Consequences (a)-(e) of those definitions. This fact will be used in the proof of Lemmas 8.7.
Let #$\#_{\epsilon}(T^{ex}_{k})$ denote the number of $\epsilon$-labelled leaves in $T^{ex}_{k}.$ Note that if #$\#_{\epsilon}(T^{ex}_{k})=0$, then $T^{ex}_{k}\in\mathcal{T}_{k}(m,n)$.
For notational convenience, let $\mathcal{T}^{ex}_{k}(m,n)$, $f^{ex}_{k}(\mathbf{x},T^{ex}_{k})$, and $g^{ex}_{k}(\mathbf{x})$ respectively denote the extended versions of $\mathcal{T}_{k}(m,n)$, $f_{k}(\mathbf{x},T_{k})$, and $g_{k}(\mathbf{x}).$
Lemma 8.2.
For $k\in\{1,2,\dots,m-1\}$, (a) and (b) below always hold:
  • (a)
    There exists a function $T^{\prime}_{0}\,:\,\mathcal{T}^{ex}_{k}(m,n)\rightarrow\mathcal{T}^{ex}_{0}(m,n)$ satisfying
    $\ell\left(T^{\prime}_{0}\left[T^{ex}_{k}\right]\right)=\ell\left(T^{ex}_{k}\right)\quad\mbox{and}\quad\forall j\in[m],\,q_{j}\left(T^{\prime}_{0}\left[T^{ex}_{k}\right]\right)=q_{j}\left(T^{ex}_{k}\right)$
  • (b)
    There exists a function $T^{\prime}_{k}\,:\,\mathcal{T}^{ex}_{0}(m,n)\rightarrow\mathcal{T}^{ex}_{k}(m,n)$ satisfying
    $\ell\left(T^{\prime}_{k}\left[T^{ex}_{0}\right]\right)=\ell\left(T^{ex}_{0}\right)+1\quad\mbox{and}\quad\forall j\in[m],\,q_{j}\left(T^{\prime}_{k}\left[T^{ex}_{0}\right]\right)=q_{j}\left(T^{ex}_{0}\right)$
Proof.
(a) follows directly from the fact that, from Definition 6.2, $\mathcal{T}_{k}(m,n)\subseteq\mathcal{T}_{0}(m,n)$ so $\mathcal{T}^{ex}_{k}(m,n)\subseteq\mathcal{T}^{ex}_{0}(m,n)$. Thus, simply setting $T^{\prime}_{0}\left[T^{ex}_{t}\right]=T^{ex}_{k}$ satisfes the required conditions.
Figure 10: Illustration of Lemma 8.2 case (b), describing $T^{\prime}_{k}\left[T^{ex}_{0}\right].$ Note that in $T^{\prime}_{k}\left[T^{ex}_{0}\right],$ leaf $10^{k}$ 1 is labelled with an $\epsilon$ so the tree is in $\mathcal{T}^{ex}_{k}(m,n)$ but not in $\mathcal{T}_{k}(m,n)$.
To see (b), given $T^{ex}_{0}$ consider the tree $T\in\mathcal{T}^{ex}_{0}(m,n)$ whose root is complete, with the left subtree of the root being a chain of $k-1$ intermediate-00 nodes, followed by one intermediate-$1$ node, and a leaf node assigned to $\epsilon$, and whose right subtree is $T^{ex}_{0}$. See Figure 10. Setting $T^{\prime}_{k}\left[T^{ex}_{0}\right]=T$ satisfies the required conditions. ∎
This permits proving:
Lemma 8.3.
Let $\mathbf{x}=(x_{1},\ldots,x_{m-1})\in\mathcal{R}^{m-1}.$ Then $\forall k\in\{1,\dots,m-1\},$
[…]
Proof.
Lemma 8.2 (a) implies that for all $T^{ex}_{k}\in\mathcal{T}^{ex}_{k}(m,n)$,
$g^{ex}_{0}(\mathbf{x})$$\leq$ $f^{ex}_{0}\left(\mathbf{x},T^{\prime}_{0}\left[T^{ex}_{k}\right]\right)$
$=$ $\ell\left(T^{\prime}_{0}\left[T^{ex}_{k}\right]\right)+\textstyle\sum_{j=1}^{m-1}q_{j}\left(T^{\prime}_{0}\left[T^{ex}_{k}\right]\right)\cdot x_{j}$
$=$ $\ell\left(T^{ex}_{k}\right)+\textstyle\sum_{j=1}^{m-1}q_{j}\left(T^{ex}_{k}\right)\cdot x_{j}$
$=$ $f^{ex}_{k}(\mathbf{x},T^{ex}_{k})+x_{k}.$
Because this is true for all $T^{ex}_{k}\in\mathcal{T}^{ex}_{k}(m,n)$, it immediately implies (i).
Similarly, Lemma 8.2(b) implies that for all $T^{ex}_{0}\in\mathcal{T}^{ex}_{0}(m,n)$,
$g^{ex}_{k}(\mathbf{x})+x_{k}$$\leq$ $f^{ex}_{k}\left(\mathbf{x},T^{\prime}_{k}\left[T^{ex}_{0}\right]\right)+x_{k}$
$=$ $\ell\left(T^{\prime}_{t}\left[T^{ex}_{0}\right]\right)+\textstyle\sum_{j=1}^{m-1}q_{j}\left(T^{\prime}_{t}\left[T^{ex}_{0}\right]\right)\cdot x_{j}$
$=$ $\ell\left(T^{ex}_{0}\right)+\textstyle\sum_{j=1}^{m-1}q_{j}\left(T^{ex}_{0}\right)\cdot x_{j}+1$
$=$ $f^{ex}_{0}(\mathbf{x},T^{ex}_{0})+1.$
Because this is true for all $T^{ex}_{0}\in\mathcal{T}^{ex}_{0}(m,n)$, it immediately implies (ii). ∎
Plugging $x_{k}=0$ into (i) and $x_{k}=1$ into (ii) proves
Corollary 8.4.
Let $\mathbf{x}\in[0,1]^{m-1}$ and $k\in\{1,\ldots,m-1\}.$Then
  • If $x_{k}=0,$ $g^{ex}_{0}(\mathbf{x})-g^{ex}_{k}(\mathbf{x})\leq 0.$
  • If $x_{k}=1,$ $g^{ex}_{k}(\mathbf{x})-g^{ex}_{0}(\mathbf{x})\leq 0.$
Note that this is exactly Lemma 7.2 but written for extended binary AIFV-$m$ coding trees rather than normal ones.
While it is not necessarily true that $\forall\mathbf{x}\in\mathcal{R}^{m-1},$ $g^{ex}_{k}(\mathbf{x})=g_{k}(\mathbf{x})$, we can prove that if $n$ is large enough, they coincide in the unit hypercube.
Lemma 8.5.
Let $n\geq 2^{m-1}.$ Then, for all $k\in[m]$ and $\forall\mathbf{x}\in[0,1]^{m-1}$, $g^{ex}_{k}(\mathbf{x})=g_{k}(\mathbf{x}).$
Plugging this Lemma into Corollary 8.4 immediately proves Lemma 7.2. It therefore only remains to prove the correctness of Lemma 8.5 .
8.1 Proving Lemma 8.5
The proof of Lemma 8.5 is split into two parts, The first justifies simplifying the structure of AIFV-$m$ trees. The second uses these properties to actually prove Lemma 8.5.
8.1.1 Further Properties of minimum-cost $T_{k}$ trees
Definitions 6.1 and 6.2 are very loose and technically permit many scenarios, e.g., the existence of more than one intermediate-$1$ node in a $T_{k}$ tree or a chain of intermediate-00 nodes descending from the root of a $T_{0}$ tree. These scenarios will not actually occur in trees appearing in minimum-cost codes. The next lemma lists some of these scenarios and justifies ignoring them. This will be needed in the actual proof of Lemma 8.5 in Section 8.1.2
Definition 8.6.
A node $v$ is a left node of $T_{k},$ if $v$ corresponds to codeword $0^{t}$ for some $t\geq 0$.
Note that for $k>0,$ $T_{k}$ contains exactly $k+1$ left nodes. With the exception of $0^{k},$ which must be an intermediate-$1$ node, the other left nodes can technically be any of complete, intermediate-00 or master nodes. By definition, they cannot be intermediate-$1$ nodes.
Lemma 8.7.
Let $k\in[m]$ and $T^{\prime}_{k}\in\mathcal{T}^{ex}_{k}(m,n).$ Then there exists $T_{k}\in\mathcal{T}^{ex}_{k}(m,n)$ satisfying
#$\#_{\epsilon}(T_{k})=\#_{\epsilon}(T^{\prime}_{k}),\quad\ell(T_{k})\leq\ell(T^{\prime}_{k})\quad\mbox{and}\quad\forall j\in[m],\,q_{j}(T_{k})=q_{j}(T^{\prime}_{k})$ (13)
and the following five conditions:
  • (a)
    The root of $T_{k}$ is not an intermediate-$1$ node;
  • (b)
    The root of $T_{0}$ is not an intermediate-00 node;
  • (c)
    If $v$ is an intermediate-$1$ node in $T_{k},$ then the parent of $v$ is not an intermediate-$1$ node;
  • (d)
    If $v$ is a non-root intermediate-00 node in $T_{k}$ and, furthermore, if $k>0,$ then $v$ is not a left node, then the parent of $v$ is either a master node or an intermediate-00 node;
  • (e)
    If $v$ is an intermediate-$1$ node in $T_{k}$ then $k>0$ and $v=0^{k}.$
    This implies that $v=0^{k}$ is the unique intermediate-$1$ node in $T_{k}.$
Proof.
If a $T^{\prime}_{k}$ tree does not satisfy one of conditions (a)-(d), we first show that it can be replaced by a $T_{k}$ tree with one fewer nodes satisfying (13). Since this process cannot be repeated forever, a tree satisfying all of the conditions (a)-(d) and satisfying (13) must exist.
Before starting, we emphasize that none of the transformations described below adds or removes $\epsilon$-leaves, so #$\#_{\epsilon}(T_{k})=\#_{\epsilon}(T^{\prime}_{k}).$
If condition (a) is not satisfied in $T^{\prime}_{k},$ then, from Consequence (e) following Definition 6.2, $k=0$. Let $r$ be the intermediate-$1$ root of $T^{\prime}_{0}$ and $v$ its child. Create $T_{0}$ by removing $r$ and making $v$ the root. Then (13) is valid (with $\ell(T_{0})<\ell(T^{\prime}_{0})$).
If condition (b) is not satisfied, let $r$ be the intermediate-00 root of $T^{\prime}_{0}$ and $v$ its child. Create $T_{0}$ by removing $r$ and making $v$ the root. Then again (13) is valid (with $\ell(T_{0})<\ell(T^{\prime}_{0})$). Note that this argument fails for $k\not=0$ because removing the edge from $r$ to $v$ would remove the node corresponding to word $10^{k}$ 1 and $T_{k}$ would then no longer be in $\mathcal{T}_{k}(m,n).$
If condition (c) is not satisfied in $T^{\prime}_{k},$ let $u$ be a intermediate-$1$ node in $T^{\prime}_{k}$ whose child $v$ is also an intermediate-$1$ node. Let $w$ be the unique child of $v.$ Now create $T_{k}$ from $T^{\prime}_{k}$ by pointing the $1$-edge leaving $u$ to $w$ instead of $v,$ i.e., removing $v$ from the tree. Then $\ell(T_{k})\leq\ell(T^{\prime}_{k})$ and (13) is valid.
Figure 11: Illustration of transformations in cases (c) and (d) of Lemma 8.7. Note that the illustration of the second subcase of (d) assumes that $v$ is the $1$-child of $u$. The case in which $v$ is the 00-child is symmetric.
If condition (d) is not satisfied in $T^{\prime}_{k},$ let $v$ be an intermediate-00 node in $T^{\prime}_{k}$ and suppose that its parent $u$ is either an intermediate-$1$ node or a complete node. Let $w$ be the unique child of $v.$ Now create $T_{k}$ from $T^{\prime}_{k}$ by taking the pointer from $u$ that was pointing to $v$ and pointing it to $w$ instead, i.e., again removing $v$ from $T^{\prime}_{k}$. Again, $\ell(T_{k})\leq\ell(T^{\prime}_{k})$ and (13) is valid. Note that the condition that “if $k>0,$ then $v$ is not a left node”, ensures that Condition 2 from Definition 6.2 is not violated.
We have shown that for any $T^{\prime}_{k},$ there exists a $T_{k}$ satisfying (a)-(d) and (13).
Now assume that conditions (a)-(d) are satisfied in $T^{\prime}_{k},$ but condition (e) is not. From (a), $v$ is not the root of $T_{k}$ so $u,$ the parent of $v,$ exists. From (c), $u$ is not an intermediate-$1$ node and from the definition of master nodes $u$ is not a master node. Thus $u$ is either a complete node or an intermediate-00 node. Let $w$ be the unique ($1$-child) of $v.$ From (c), $w$ cannot be an intermediate-$1$ node; from (d), $w$ cannot be an intermediate-00 node. So $w$ is either a complete or master node.
Now create $T_{k}$ from $T^{\prime}_{k}$ by taking the pointer from $u$ that was pointing to $v$ and pointing it to $w$ instead, i.e., again removing $v$ from $T^{\prime}_{k}.$
Note that after the transformation, it is easy to see that $\ell(T_{k})\leq\ell(T^{\prime}_{k})$ and (13) is valid. Note that since $w$ is either a complete or master node, pointing $u$ to $w$ does not affect the master nodes above $v.$ It only remains to show that this pointer redirection is a permissible operation on $T_{k}$ trees, i.e., that $T_{k}$ does not violate Condition 2 from Definition 6.2.
Since (e) is not satisfied, $v\neq 0^{k}.$ There are two cases:
Case 1: $k>0$ and $v=0^{r}.$
This is not possible for $r>k$ because $0^{r}\not\in T_{k}$. It is also not possible for $0<r<k,$ because in that case, $0^{r}$ has a 00-child so it can not be an intermediate-$1$ node.
Case 2: $k=0$ or $v\not=0^{r}$ for any $r>0.$
If $k=0$ then, trivially, Condition 2 from Definition 6.2 cannot be violated. If $k\not=0,$ then the transformation described leaves $0^{k}$ as an intermediate-$1$ node so Condition 2 from Definition 6.2 is still not violated.
Figure 12: Illustration of transformation in case (e) of Lemma 8.7. $v$ is always an intermediate-$1$ node and $w$ can be either a master node or a complete node. $z$ is a master node of degree $t$. In the case in which $u$ is a complete node, $v$ may be either the left or right child of $u.$ Both cases are illustrated.
This operation of removing an intermediate-$1$ node can be repeated until condition (e) is satisfied. ∎
This lemma has two simple corollaries.
Corollary 8.8.
There exists a minimum-cost AIFV-$m$ code $\mathbf{T}=(T_{0},\ldots,T_{m-1})$ such that $\forall k\in[m],$ tree $T_{k}$ satisfies conditions (a)-(e) of Lemma 8.7.
Proof.
Let $\mathbf{T}^{\prime}=(T^{\prime}_{0},\ldots,T^{\prime}_{m-1})$ be a minimum cost AIFV-$m$ code. For each $k\in[m],$ let $T_{k}$ be the tree satisfying conditions (a)-(e) and equation (13) and set $\mathbf{T}=(T_{0},\ldots,T_{m-1}).$ Since $\mathbf{Q}(\mathbf{T})=\mathbf{Q}(\mathbf{T}^{\prime}),$ $\bm{\pi}(\mathbf{T})=\bm{\pi}(\mathbf{T}^{\prime}).$ Since $\forall k,$ $\ell(T_{k})\leq\ell(T^{\prime}_{k})$,
${\rm cost}(\mathbf{T})=\sum_{k=0}^{m-1}\ell(T_{k})\cdot\pi_{k}(\mathbf{T})=\sum_{k=0}^{m-1}\ell(T_{k})\cdot\pi_{k}(\mathbf{T}^{\prime})\leq\sum_{k=0}^{m-1}\ell(T^{\prime}_{k})\cdot\pi_{k}(\mathbf{T}^{\prime})={\rm cost}(\mathbf{T}^{\prime}).$
But $\mathbf{T}^{\prime}$ was a minimum-cost AIFV-$m$ code so $\mathbf{T}$ must be one as well. ∎
The corollary implies that, our algorithmic procedures, for all $k\in[m],$ may assume that all $T_{k}\in\mathcal{T}_{k}(m,n)$ satisfy conditions (a)-(e) of Lemma 8.7.
This assumption permits bounding the height of all $T_{k}$ trees. The following corollary was needed in the proof of Lemma 7.1.
Corollary 8.9.
For all $T_{k}\in\mathcal{T}_{k}(m,n),$ the height of $T_{k}$ is at most $n(m+1).$
Proof.
Let $T_{k}\in\mathcal{T}_{k}(m,n)$ satisfy conditions (a)-(e) of Lemma 8.7.
  • Let $\ell\geq 1$ be the number of leaves in $T_{k}.$ Since all leaves are master nodes, $T_{k}$ contains $n-\ell$ non-leaf master nodes.
  • Every complete node in $T_{k}$ must contain at least one leaf in each of its left and right subtrees, so the number of complete nodes in $T_{k}$ is at most $\ell-1.$
  • $T_{k}$ contains no intermediate-$1$ node if $k=0$ and one intermediate-$1$ node if $k>1.$
  • If $k=0,$ each intermediate-00 node in $T_{k}$ can be written as $v0^{t}$ for some non-leaf master node $v$ and $t\leq m.$ If $k>0,$ all intermediate-00 nodes in $T_{k}$ with the possible exceptions of the left nodes $0^{r},$ $r\leq k,$ can be written as $v0^{t}$ for some non-leaf master node $v$ and $t\leq m.$
    So, the total number of intermediate-00 nodes in the tree is at most $(n-\ell)m+(k+1).$
  • The total number of non-leaf nodes in the tree is then at most
    $(n-\ell)+(\ell-1)+1+(n-\ell)m+k+1$ $=$ $n+(n-\ell)m+(k+1)$
    $\leq$ $n+(n-1)m+m$
    $=$ $n(m+1).$
Thus any path from a leaf of $T_{k}$ to its root has length at most $n(m+1).$
We note that this bound is almost tight. Consider a $T_{0}$ tree which has only one leaf and with all of the other master nodes (including the root) being master nodes of degree $m.$ This tree is just a chain from the root to the unique leaf, and has of length $(n-1)(m+1).$
8.1.2 The Actual Proof of Lemma 8.5
It now remains to prove Lemma 8.5, i.e., that if $n\geq 2^{m-1},$ then, for all $k\in[m]$ and $\forall\mathbf{x}\in[0,1]^{m-1}$, $g^{ex}_{k}(\mathbf{x})=g_{k}(\mathbf{x}).$
Proof.
(of Lemma 8.5.)
Recall that $\mathcal{T}_{k}(m,n)\subseteq\mathcal{T}^{ex}_{k}(m,n)$ so, $\forall\mathbf{x}\in\mathcal{R}^{m-1},$ $g^{ex}_{k}(\mathbf{x})\leq g_{k}(\mathbf{x}).$
Fix $\mathbf{x}\in[0,1]^{m-1}$. Now let $T^{ex}_{k}\in\mathcal{T}^{ex}_{k}(m,n)$ be a code tree satisfying
  • (i)
    $f^{ex}_{k}(\mathbf{x},T^{ex}_{k})=\min_{T\in\mathcal{T}^{ex}_{k}(m,n)}f^{ex}_{k}(\mathbf{x},T)=g^{ex}_{k}(\mathbf{x})$ and
  • (ii)
    among all trees satisfying (i), $T^{ex}_{k}$ is a tree minimizing the number of leaves assigned an $\epsilon$.
Because (13) in Lemma 8.7 keeps $q_{j}(T^{ex}_{k})$ the same and cannot increase $\ell(T^{ex}_{k}),$ we may also assume that $T^{ex}_{k}$ satisfies conditions (a)-(e) of Lemma 8.7.
Since $f^{ex}_{k}(\mathbf{x},T^{ex}_{k})=g^{ex}_{k}(\mathbf{x})$, to prove the lemma, it thus suffices to show that $T^{ex}_{k}\in\mathcal{T}_{k}(m,n).$ This implies that $g_{k}(\mathbf{x})\leq g^{ex}_{k}(\mathbf{x})$ so $g_{k}(\mathbf{x})=g^{ex}_{k}(\mathbf{x}).$
Suppose to the contrary that $T^{ex}_{k}\not\in\mathcal{T}_{k}(m,n)$, i.e., $T^{ex}_{k}$ contains a leaf $l_{\epsilon}$ assigned an $\epsilon$. Let $a_{\epsilon}$ denote the closest (lowest) non-intermediate-00 ancestor of $l_{\epsilon}$ and $b_{\epsilon}$ be the child of $a_{\epsilon}$ that is the root of the subtree containing $l_{\epsilon}.$
By the definition of $a_{\epsilon},$ either $l_{\epsilon}=b_{\epsilon}$ or $l_{\epsilon}=b_{\epsilon}0^{t}$ where $b_{\epsilon}0^{r},$ $0\leq r<t$ are all intermediate-00 nodes for some $t>0.$ Note that if $k>0,$ no left node is a leaf so, in particular neither $l_{\epsilon}$ or $b_{\epsilon}$ can be a left node.
We can therefore apply Lemma 8.7 (d) to deduce that, if $b_{\epsilon}$ was an intermediate-00 node, $a_{\epsilon}$ must be a master node. Thus, if $a_{\epsilon}$ is a complete node or intermediate-$1$ node, $b_{\epsilon}=l_{\epsilon}.$ If $a_{\epsilon}$ is a master node, then $b_{\epsilon}=a_{\epsilon}0$ and $l_{\epsilon}=a_{\epsilon}0^{t+1}$ so $a_{\epsilon}$ is a master node of degree $t.$
Now work through the three cases:
  • (a)
    $a_{\epsilon}$ is a complete node and $b_{\epsilon}=l_{\epsilon}:$ (See Figure 13 (a))
    Let $c_{\epsilon}$ be the other child of $a_{\epsilon}$ (that is not $b_{\epsilon}$). Now remove $b_{\epsilon}$. If $c_{\epsilon}$ was not already the $1$-child of $a_{\epsilon},$ make $c_{\epsilon}$ the $1$-child of $a_{\epsilon}.$
    Figure 13: Illustration of the first two cases of the proof of Lemma 8.5. In case (a) $a_{\epsilon}$ is complete and $c_{\epsilon}$ can be its $1$-child or its 00-child. $T^{\prime}$ is the subtree rooted at $c_{\epsilon}.$ In case (b) $a_{\epsilon}$ is a master node and $\ell_{\epsilon}$ is connected to it via a chain of intermediate-00 nodes.
    The above transformation makes $a_{e}$ an intermediate-$1$ node. The resulting tree remains a valid tree in $\mathcal{T}^{ex}_{k}(m,n)$ preserving the same cost $f^{ex}_{k}(\mathbf{x},T^{ex}_{k})$ but reducing the number of leaves assigned to $\epsilon$ by $1$. This contradicts the minimality of $T^{ex}_{k}$, so this case is not possible.
  • (b)
    $a_{\epsilon}$ is a master node of degree $t>0:$ (See Figure 13 (b))
    Let $\sigma_{i}$ be the source symbol assigned to $a_{\epsilon}.$ Next remove the path from $a_{\epsilon}$ to $l_{\epsilon},$ converting $a_{\epsilon}$ to a leaf, i.e., a master node of degree 00. This reduces [^9] This is the only location in the proof that uses $\mathbf{x}\in[0,1]^{m-1}.$ $f^{ex}_{k}(\mathbf{x},T^{ex}_{k})$ by $p_{i}\cdot x_{t}\geq 0$ and the number of leaves assigned to $\epsilon$ by $1$. The resulting code tree is still in $\mathcal{T}^{ex}_{k}(m,n)$ and contradicts the minimality of $T^{ex}_{k}$, so this case is also not possible.
    Figure 14: Illustration of case (c) of the proof of Lemma 8.5. $a_{\epsilon}=0^{k}$ is an intermediate-$1$ node and $\beta_{\epsilon}=0^{k}1.$ From definition 6.2 (2), this $b_{\epsilon}$ must exist in every $T_{k},$ so it may not be removed. $v_{i}$ is a deepest leaf in $T_{k},$ which is “swapped” with $b_{\epsilon}$.
  • (c)
    $a_{\epsilon}$ is an intermediate-$1$ node and $b_{\epsilon}=l_{\epsilon}:$ (See Figure 14)
    From Lemma 8.7 (a) and (e), $k\not=0$ and $a_{\epsilon}=0^{k}$, $b_{\epsilon}=0^{k}1$. So, $b_{\epsilon}$ is at depth $k+1$.
    Since cases (a) and (b) cannot occur, $10^{k}$ 1 is the unique leaf assigned an $\epsilon.$ All other master nodes must be assigned some source symbol.
    Let $\sigma_{i}$ be the deepest source symbol in the tree that is assigned to a non-left node and $v_{i}$ be the master node to which $\sigma_{i}$ is assigned. Since $n\geq 2^{m}\geq k,$ such a $v_{i}$ must exist.
    Since every master node in $T^{ex}_{k}$ except for $10^{k}$ 1 is assigned a source symbol, from consequence (a) following Definition 6.2, $v_{i}$ is also a leaf.
    Now swap $b_{\epsilon}$ and $v_{i}.$ i.e., assign $\sigma_{i}$ to $10^{k}$ 1 and $\epsilon$ to the node that used to be $\sigma_{i}.$
    The resulting tree is still in $\mathcal{T}^{ex}_{k}(m,n).$ Furthermore, since the degree of all master nodes associated with source symbols remains unchanged, $\sum_{j=1}^{m-1}q_{j}(T^{ex}_{k})\cdot x_{j}$ remains unchanged.
    Now consider $\ell\left(T^{ex}_{k},\sigma_{i}\right).$
    If $\ell\left(T^{ex}_{k},\sigma_{i}\right)>k+1$ before the swap, then $\ell\left(T^{ex}_{k}\right)$, and therefore $f^{ex}_{k}(\mathbf{x},T^{ex}_{k}),$ are decreased by at least $p_{i}>0$ by the swap. This contradicts the minimality of $T^{ex}_{k}$, so this is not possible.
    If $\ell\left(T^{ex}_{k},\sigma_{i}\right)=k+1$ before the swap, then $\ell\left(T^{ex}_{k}\right),$ and therefore $f^{ex}_{k}(\mathbf{x},T^{ex}_{k}),$ remain unchanged. Thus the new tree satisfies conditions (i) and (ii) at the beginning of the proof. Since the original $T^{ex}_{k}$ satisfied conditions (a)-(e) of Lemma 8.7, the new tree does as well. Because of the swap, the new tree does not satisfy case (c), so it must be in case (a) or (b). But, as already seen, neither of those cases can occur. So, $\ell\left(T^{ex}_{k},\sigma_{i}\right)\not=k+1.$
We have therefore proven that, if $T^{ex}_{k}\not\in\mathcal{T}_{k}(m,n)$, then it must satisfy case (c) with $\ell\left(T^{ex}_{k},\sigma_{i}\right)<k+1.$ We now claim that if $n\geq 2^{k}$, then $\ell\left(T^{ex}_{k},\sigma_{i}\right)\geq k+1.$ This, combined with $k\leq m-1,$ would prove the lemma.
To prove the claim, let $T^{ex}_{k}\in\cup_{n\geq 1}\mathcal{T}^{ex}_{k}(m,n)$ be some tree that maximizes the number of master nodes at depths $\leq k.$ Call such a tree a $k$-maximizing tree.
Suppose $T^{ex}_{k}$ contained some non-leaf master node $v$ of depth $<k$. Transform $v$ into a complete node by giving it a $1$-child that is a leaf. The resulting tree has the same number of master nodes, so it is also a $k$-maximizing tree, but contains fewer non-leaf master nodes. Repeating this operation yields a $k$-maximizing tree in which all of the master nodes at depth $<k$ are leaves.
Next note that a $k$-maximizing tree cannot contain any leaves of depth $<k$ because, if it did, those leaves could be changed into internal nodes with two leaf children, contradicting the definition of a $k$-maximizing tree.
We have thus seen that there is a $k$-maximizing tree $T^{ex}_{k}$ in which all of the master nodes are on level $k$. But, any binary tree can have at most $2^{k}$ total nodes on level $k$ and $0^{k}$ is not a master node in a type-$k$ tree, so $T^{ex}_{k}$ contains at most $12^{k}-$ 1 master nodes. This implies that any type-$k$ tree with $\geq 2^{k}$ master nodes must have some (non-left) master node at depth $\geq k+1$, proving the claim and thus the lemma. ∎
9 Conclusion and Directions for Further Work
The first part of this paper introduced the minimum-cost Markov chain problem. We then showed how to translate it into the problem of finding the highest point in the Markov Chain Polytope $\mathbb{H}.$ In particular, Lemma 4.6 in Section 4.3 identified the problem specific information that is needed to use the Ellipsoid algorithm to solve the problem in polynomial time.
This was written in a very general form so that it could be applied to solve problems other than binary Almost Instantaneous Fixed-to-Variable-$m$ (AIFV-$m$) coding. For example, recent work [2] uses this Lemma to derive polynomial time algorithms for AIVF-coding, a generalization of Tunstall coding which previously [17, 18] could only be solved in exponential time using an iterative algorithm.
Another possible problem use of Lemma 4.6 would be the construction of optimal codes for finite-state noiseless channels. [18] recently showed how to frame this problem as a minimum cost Markov chain one and solve it using the iterative algorithm. This problem would definitely fit into the framework of Lemma 4.6. Unfortunately, the calculation of the corresponding $\mathbf{S}(\mathbf{x})$ in [18], needed by Lemma 4.6 as a problem-specific separation oracle, is done using Integer Linear Programming and therefore requires exponential time. The development of a polynomial time algorithm for calculating the $\mathbf{S}(\mathbf{x})$ would, by Lemma 4.6, immediately yield polynomial time algorithms for solving the full problem.
The second part of the paper restricts itself to binary AIFV-$m$ codes, which were the original motivation for the minimum-cost Markov Chain problem. These are $m$-tuples of coding trees for lossless-coding that can be modelled as a Markov chain. We derived properties of AIFV-$m$ coding trees that then permitted applying Lemma 4.6. This yielded the first (weakly) polynomial time algorithm for constructing minimum cost binary AIFV-$m$ codes.
There are still many related open problems to resolve. The first is to note that our algorithm is only weakly polynomial, since its running time is dependent upon the actual sizes needed to encode the transition probabilities of the Markov Chain states in binary. For example, in AIFV-$m$ coding, this is polynomial in the number of bits needed to encode the probabilities of the words in the source alphabet. An obvious goal would be, at least for AIFV-$m$ coding, to find a strongly polynomial time algorithm, one whose running time only depends upon $n.$
A second concerns the definition of permissible Markov chains. Definition 2.1 requires that $\forall k\in[m],\,\forall S_{k}\in{\mathbb{S}}_{k},$ $q_{0}(S_{k})>0.$ This guarantees that every permissible Markov Chain $\mathbf{S}$ has a unique stationary distribution. It is also needed as a requirement for Theorem 1 in [1], which is used to guarantee that there exists a distinctly-typed intersection point in $\mathbb{H}$ (Corollary 3.2 only guarantees that every distinctly-typed intersection point is on or above $\mathbb{H}$). The open question is whether $q_{0}(S_{k})>0$ is actually needed or whether the looser requirement that every $\mathbf{S}$ has a unique stationary distribution would suffice to guarantee that $\mathbb{H}$ contains some distinctly-typed intersection point.
A final question would return to the iterative algorithm approach of [23, 22, 4, 5]. Our new polynomial time algorithm is primarily of theoretical interest and, like most Ellipsoid based algorithms, would be difficult to implement in practice. Perhaps the new geometric understanding of the problem developed here could improve the performance and analysis of the iterative algorithms.
As an example, the iterative algorithm of [4, 5, 22, 23] can now be interpreted as moving from point to point in the set of distinctly-typed intersection points (of associated hyperplanes), never increasing the cost of the associated Markov chain, finally terminating at the lowest point in this set.
This immediately leads to a better understanding of one of the issues with the iterative algorithms for the AIFV-$m$ problem.
As noted in Section 2.3, the algorithm must solve for $\mathbf{S}(\mathbf{x})$ at every step of the algorithm. As noted in Section 4, this can be done in polynomial time if $\mathbf{x}\in[0,1]^{m-1}$ but requires exponential time integer linear programming if $\mathbf{x}\not\in[0,1]^{m-1}.$ A difficulty with the iterative algorithm was that it was not able to guarantee that at every step, or even at the final solution, $\mathbf{x}\in[0,1]^{m-1}.$ With our new better understanding of the geometry of the Markov Chain polytope for the AIFV-$m$ problem, it might now be possible to prove that the condition $\mathbf{x}\in[0,1]^{m-1}$ always holds during the algorithm or develop a modified iterative algorithm in which the condition always holds.
Acknowledgement: The authors would like to thank Reza Hosseini Dolatabadi and Arian Zamani for the generation of Figures 2 and 3.